### v1

# Non-standard analysis for coherent risk estimation: hyperfinite representations, discrete Kusuoka formulae, and plug-in asymptotics

Abstract.
We develop a non-standard analysis framework for coherent risk measures and their finite-sample analogues, coherent risk estimators, building on recent work of Aichele, Cialenco, Jelito, and Pitera. Coherent risk measures on $L^{\infty}$ are realised as standard parts of internal support functionals on Loeb probability spaces, and coherent risk estimators arise as finite-grid restrictions.
Our main results are: (i) a hyperfinite robust representation theorem that yields, as finite shadows, the robust representation results for coherent risk estimators; (ii) a discrete Kusuoka representation for law-invariant coherent risk estimators as suprema of mixtures of discrete expected shortfalls on $\{k/n:k=1,\ldots,n\}$; (iii) uniform almost sure consistency (with an explicit rate) for canonical spectral plug-in estimators over Lipschitz spectral classes; (iv) a Kusuoka-type plug-in consistency theorem under tightness and uniform estimation assumptions; (v) bootstrap validity for spectral plug-in estimators via an NSA reformulation of the functional delta method (under standard smoothness assumptions on $F_{X}$); and (vi) asymptotic normality obtained through a hyperfinite central limit theorem.
The hyperfinite viewpoint provides a transparent probability-to-statistics dictionary: applying a risk measure to a law corresponds to evaluating an internal functional on a hyperfinite empirical measure and taking the standard part. We include a standard self-contained introduction to the required non-standard tools.
Key words and phrases:
coherent risk measure, coherent risk estimator, spectral risk, Kusuoka representation, non-standard analysis, Loeb measure, order statistics, L-statistics, hyperfinite probability
2020 Mathematics Subject Classification:
60B10, 60F15, 62G30, 91G70, 03H05
IM CAS (RVO 67985840).
1. Introduction
1.1. Background and motivation
The measurement of financial risk occupies a central position in modern quantitative finance, insurance, and regulatory frameworks. Since the seminal work of Artzner, Delbaen, Eber, and Heath [2], the axiomatic approach to risk measures has provided both theoretical foundations and practical guidance for risk management. A coherent risk measure (CRM) is a functional $\rho$ acting on random variables (representing profit and loss distributions) that satisfies four economically motivated axioms: monotonicity, cash additivity, positive homogeneity, and subadditivity. (Throughout, we adopt the P&L sign convention: $X$ represents profit, so losses are negative and $\rho(X)$ evaluates required capital; hence the (X)(-X) in dual representations.)
The fundamental structural result for CRMs is the robust representation theorem, which asserts that under mild regularity conditions, any coherent risk measure can be written as a supremum of expectations under a family of probability measures:
$\rho(X)=\sup_{Q\in\mathcal{Q}}\mathsf{E}_{Q}[-X].$
This representation reveals the dual nature of coherent risk: the set $\mathcal{Q}$ encodes model uncertainty or stress scenarios, whilst the supremum captures a worst-case perspective essential to prudent risk management.
Whilst the theory of CRMs is well-developed at the population level—where one has access to the full distribution of XX—practical applications invariably require estimation from finite samples. This raises the statistical question: what is the appropriate analogue of coherent risk when one observes only a sample $(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$?
This question was recently addressed by Aichele, Cialenco, Jelito, and Pitera [1], who introduced the concept of a coherent risk estimator (CRE): a functional $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ satisfying the same four coherence axioms, now interpreted coordinatewise on the sample space. Their work establishes that CREs admit finite-dimensional robust representations as suprema of linear functionals, with law-invariant and comonotonic variants corresponding to L-estimators based on order statistics.
1.2. The non-standard analysis perspective
The purpose of this paper is to develop a unified framework for CRMs and CREs using non-standard analysis (NSA). This approach, originating in the work of Abraham Robinson [16], extends the real numbers to the hyperreals $^{*}\mathbb{R}$, which include infinitesimal and infinite elements. Whilst initially developed for logical and foundational purposes, NSA has proven to be a powerful tool in probability theory, particularly through the Loeb measure construction [15], which produces genuine $\sigma$-additive probability spaces from internal hyperfinite structures.
From the NSA perspective, the objects of interest have particularly transparent interpretations:
  • Hyperfinite probability spaces. An internal hyperfinite set $I_{N}=\{1,2,\ldots,N\}$ with $N\in{}^{*}\mathbb{N}$ infinite, equipped with the counting probability measure $\mu_{N}(A)=|A|/N$, becomes a genuine probability space via the Loeb construction. This provides a “bridge” between discrete and continuous probability.
  • Coherent risk measures as standard parts. A CRM $\rho$ on $L^{\infty}$ can be expressed as the standard part of a hyperfinite support functional: dual measures $Q$ correspond to hyperfinite weight vectors $a=(a_{1},\ldots,a_{N})$ with $\sum_{k=1}^{N}a_{k}=1$, and expectations under $Q$ become hyperfinite sums.
  • CREs as finite shadows. By taking $N=n$ finite, the hyperfinite representation specialises to the finite-dimensional robust representation of CREs. The passage from CRMs to CREs is simply the restriction of the hyperfinite picture to a finite grid.
  • Plug-in estimators as hyperfinite L-statistics. On an infinite hyperfinite i.i.d. sample $(X_{1},\ldots,X_{N})$, spectral risk measures become standard parts of hyperfinite L-statistics. Consistency results emerge from the hyperfinite law of large numbers and quantile convergence.
This unified viewpoint offers several advantages. First, it provides a single conceptual framework encompassing both the population-level theory of CRMs and the finite-sample theory of CREs. Second, it simplifies certain asymptotic arguments by allowing one to work internally with the entire probability space at once, taking standard parts only at the end. Third, it suggests natural generalisations and new results, including uniform consistency over families of risk measures and bootstrap validity.
1.3. Main contributions
The principal contributions of this paper are as follows.
  1. (1)
    Hyperfinite robust representation (Section 4). We establish that coherent risk measures on $L^{\infty}$ are standard parts of hyperfinite support functionals, providing a unified proof of the finite-sample robust representation theorems for CREs (Theorems 2.9, 2.10, and 2.11).
  2. (2)
    Discrete Kusuoka representation (Section 5). We prove that every law-invariant CRE admits a representation as a supremum over mixtures of discrete expected shortfalls at grid points (Theorem 5.5). This is the finite-sample analogue of Kusuoka’s celebrated representation theorem for law-invariant CRMs on atomless spaces.
  3. (3)
    Uniform spectral consistency (Section 7). We establish uniform almost sure consistency for spectral plug-in CREs over Lipschitz families of spectra, with an explicit rate of convergence (Theorem 7.6 and Corollary 7.8).
  4. (4)
    Kusuoka plug-in consistency (Section 8). We prove a general consistency theorem for Kusuoka-type plug-in estimators under tightness and uniform estimation conditions (Theorem 8.1).
  5. (5)
    Hyperfinite bootstrap validity (Section 9). We show that the internal resampling scheme yields the correct Gaussian limit, via an NSA reformulation of the bootstrap delta method (Theorem 9.4).
  6. (6)
    Asymptotic normality (Section 10). We derive the asymptotic distribution of spectral plug-in estimators via the hyperfinite central limit theorem (Theorem 10.5).
1.4. Organisation of the paper
Section 2 recalls the definitions of coherent risk measures and estimators, following [1] and [11]. Section 3 provides a self-contained introduction to non-standard analysis, covering hyperreals, hyperfinite sets, and Loeb measures at the level required for our applications. Section 4 develops the hyperfinite representation of CRMs and uses it to derive the finite-sample representation theorems for CREs. Section 5 establishes the discrete Kusuoka representation. Section 6 treats spectral risk measures and their hyperfinite L-statistic representations. Section 7 proves the uniform spectral consistency theorem. Section 8 extends to general Kusuoka-type plug-in estimators. Section 9 establishes bootstrap validity. Section 10 derives asymptotic normality. Section 11 briefly discusses the extension to Orlicz hearts.
The role of NSA varies across sections: in Sections 45, NSA provides a unifying language for dual representations, showing that CREs are finite shadows of CRMs; in Sections 710, NSA serves as a probability-to-statistics transfer device, converting population-level results (Glivenko–Cantelli, CLT, bootstrap validity) to sample-level statements via the standard-part map.
1.5. Notation
Throughout, $(\Omega,\mathscr{G},\mathsf{P})$ denotes a probability space. For brevity, we write $L^{0}$ for $L^{0}(\Omega,\mathscr{G},\mathsf{P})$, the space of (equivalence classes of) measurable functions, and $L^{\infty}$ for $L^{\infty}(\Omega,\mathscr{G},\mathsf{P})$, the space of essentially bounded functions. Equalities and inequalities between random variables are understood in the $\mathsf{P}$-almost sure sense.
For $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, we write $x_{1:n}\leqslant x_{2:n}\leqslant\cdots\leqslant x_{n:n}$ for the order statistics and $s(x)=(x_{1:n},\ldots,x_{n:n})$ for the sorted sample. The standard simplex is
$\Delta_{n}:=\Big\{a\in[0,\infty)^{n}:\sum_{i=1}^{n}a_{i}=1\Big\},$
and its monotone subset is
$\Delta_{n}^{\downarrow}:=\{a\in\Delta_{n}:a_{1}\geqslant a_{2}\geqslant\cdots\geqslant a_{n}\}.$
2. Coherent risk measures and coherent risk estimators
This section recalls the basic definitions and fundamental representation theorems, following the exposition in [1] and the standard reference [11].
2.1. Coherent risk measures on $L^{\infty}$
The axiomatic approach to risk measurement begins with economically motivated desiderata for a ‘good’ risk functional. We interpret $X\in L^{\infty}$ as a random profit and loss (P&L), with positive values representing gains and negative values representing losses. The quantity $\rho(X)$ represents the capital required to make the position $X$ acceptable.
Definition 2.1 (Coherent risk measure).
A functional $\rho:L^{\infty}\to\mathbb{R}$ is a coherent risk measure (CRM) if for all $X,Y\in L^{\infty}$, $m\in\mathbb{R}$, and $\lambda\geqslant 0$:
  1. (R1)
    Monotonicity: $X\leqslant Y$ implies $\rho(X)\geqslant\rho(Y)$.
  2. (R2)
    Cash additivity: $\rho(X+m)=\rho(X)-m$.
  3. (R3)
    Positive homogeneity: $\rho(\lambda X)=\lambda\rho(X)$.
  4. (R4)
    Subadditivity: $\rho(X+Y)\leqslant\rho(X)+\rho(Y)$.
The axioms have clear economic interpretations. Monotonicity states that a position with uniformly better outcomes requires less capital. Cash additivity asserts that adding a deterministic amount $m$ to the position reduces the required capital by the same amount. Positive homogeneity implies that scaling a position scales the risk proportionally. Subadditivity captures the principle of diversification: the risk of a combined portfolio does not exceed the sum of individual risks.
The fundamental structural result for CRMs is the robust representation theorem, which characterises coherent risk measures in terms of their dual sets of probability measures. We recall one version; see [8, 11] for a comprehensive treatment.
Theorem 2.2 (Robust representation on $L^{\infty}$).
Let $\rho:L^{\infty}\to\mathbb{R}$ be a coherent risk measure satisfying the Fatou property: if $(X_{n})$ is a bounded sequence converging $\mathsf{P}$-almost surely to $X$, then $\rho(X)\leqslant\liminf_{n\to\infty}\rho(X_{n})$. Then there exists a non-empty convex set $\mathcal{Q}$ of probability measures on $(\Omega,\mathscr{G})$, each absolutely continuous with respect to $\mathsf{P}$, such that
(1) $\rho(X)=\sup_{Q\in\mathcal{Q}}\mathsf{E}_{Q}[-X],\qquad X\in L^{\infty}.$
If, in addition, the set of Radon–Nikodym derivatives $\{dQ/d\mathsf{P}:Q\in\mathcal{Q}\}\subset L^{1}$ can be chosen $\sigma(L^{1},L^{\infty})$-compact (equivalently, uniformly integrable and $\sigma(L^{1},L^{\infty})$-closed), then the supremum is attained for each $X$.
The set $\mathcal{Q}$ is often interpreted as a family of “stress scenarios” or “generalised probability assessments”, and the representation (1) expresses the risk as a worst-case expected loss over these scenarios.
2.2. Law invariance and spectral representation
A particularly important class of CRMs are those depending only on the distribution of the random variable.
Definition 2.3 (Law invariance).
A CRM $\rho$ is law-invariant if $\rho(X)=\rho(Y)$ whenever $X$ and $Y$ have the same distribution under $\mathsf{P}$.
For law-invariant CRMs on atomless probability spaces, Kusuoka [14] established a remarkable representation in terms of expected shortfall:
(2) $\operatorname{ES}_{\alpha}(X):=-\frac{1}{\alpha}\int_{0}^{\alpha}q_{X}(u)\,du,\qquad\alpha\in(0,1],$
where $q_{X}:(0,1)\to\mathbb{R}$ is the lower quantile function of $X$, $q_{X}(\alpha):=\inf\{x\in\mathbb{R}:F_{X}(x)\geqslant\alpha\}$.
Remark 2.4 (Parameterisation convention).
In our convention, the parameter $\alpha\in(0,1]$ is a tail probability: small $\alpha$ corresponds to the worst tail of the profit distribution (largest losses). This differs from the regulatory convention for loss distributions, where ES is often quoted at a “confidence level” close to $1$ (e.g., $\operatorname{ES}_{0.975}$ for the $2.5\%$ tail of losses). The two are related by $\operatorname{ES}_{\alpha}^{\text{profit}}(X)=\operatorname{ES}_{1-\alpha}^{\text{loss}}(-X)$.
Theorem 2.5 (Kusuoka representation [14]).
Let $(\Omega,\mathscr{G},\mathsf{P})$ be atomless and $\rho:L^{\infty}\to\mathbb{R}$ a law-invariant CRM satisfying the Fatou property. Then there exists a non-empty convex set $\mathcal{M}$ of probability measures on $(0,1]$ such that
(3) $\rho(X)=\sup_{\nu\in\mathcal{M}}\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha),\qquad X\in L^{\infty}.$
A special case arises when the supremum in (3) is achieved by a single measure, leading to spectral risk measures.
Definition 2.6 (Spectral risk measure).
Let $\varphi:[0,1]\to[0,\infty)$ be a function satisfying:
  1. (S1)
    $\varphi$ is non-increasing;
  2. (S2)
    $\varphi$ is bounded;
  3. (S3)
    $\int_{0}^{1}\varphi(\alpha)\,d\alpha=1$.
The spectral risk measure with spectrum $\varphi$ is
(4) $\rho_{\varphi}(X):=-\int_{0}^{1}q_{X}(\alpha)\,\varphi(\alpha)\,d\alpha.$
Spectral risk measures are coherent and law-invariant. The expected shortfall $\operatorname{ES}_{\alpha}$ corresponds to the spectrum $\varphi(u)=\alpha^{-1}\mathbf{1}_{(0,\alpha]}(u)$.
2.3. Coherent risk estimators
We now turn to the finite-sample setting. Fix $n\in\mathbb{N}$ and interpret $x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$ as a realised sample of P&L values.
Definition 2.7 (Coherent risk estimator [1, Definition 3.1]).
A mapping $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ is a coherent risk estimator (CRE) if for all $x,y\in\mathbb{R}^{n}$, $m\in\mathbb{R}$, and $\lambda\geqslant 0$:
  1. (E1)
    Monotonicity: $x_{i}\leqslant y_{i}$ for all $i$ implies $\hat{\rho}_{n}(x)\geqslant\hat{\rho}_{n}(y)$.
  2. (E2)
    Cash additivity: $\hat{\rho}_{n}(x+m\mathbf{1})=\hat{\rho}_{n}(x)-m$, where $\mathbf{1}=(1,\ldots,1)$.
  3. (E3)
    Positive homogeneity: $\hat{\rho}_{n}(\lambda x)=\lambda\hat{\rho}_{n}(x)$.
  4. (E4)
    Subadditivity: $\hat{\rho}_{n}(x+y)\leqslant\hat{\rho}_{n}(x)+\hat{\rho}_{n}(y)$.
The axioms (E1)–(E4) are the direct translations of (R1)–(R4) to the finite-sample setting. Note that we follow the sign convention of [1]: larger losses (smaller $x_{i}$) increase the required capital.
Definition 2.8 (Law invariance and comonotonicity for CREs).
A CRE $\hat{\rho}_{n}$ is law-invariant if $\hat{\rho}_{n}(x)=\hat{\rho}_{n}(\sigma(x))$ for every permutation $\sigma$ of $\{1,\ldots,n\}$. It is called comonotonic whenever $\hat{\rho}_{n}(x+y)=\hat{\rho}_{n}(x)+\hat{\rho}_{n}(y)$ whenever $x$ and $y$ are comonotonic, meaning $(x_{i}-x_{j})(y_{i}-y_{j})\geqslant 0$ for all $i,j$.
The main representation results for CREs, established in [1], are:
Theorem 2.9 (Robust representation of CREs [1, Theorem 4.1]).
A function $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ is a CRE if and only if there exists a non-empty convex set $M^{*}_{\hat{\rho}_{n}}\subseteq\Delta_{n}$ such that
(5) $\hat{\rho}_{n}(x)=\sup_{a\in M^{*}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i}),\qquad x\in\mathbb{R}^{n},$
and the supremum is attained for each $x$.
Theorem 2.10 (Law-invariant CREs [1, Theorem 4.2]).
A function $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ is a law-invariant CRE if and only if there exists a non-empty convex set $M^{s}_{\hat{\rho}_{n}}\subseteq\Delta_{n}^{\downarrow}$ such that
(6) $\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i:n}),\qquad x\in\mathbb{R}^{n},$
with the supremum attained for each $x$.
Theorem 2.11 (Comonotonic law-invariant CREs [1, Theorem 4.10]).
A function $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ is a comonotonic, law-invariant CRE if and only if there exists a unique $a\in\Delta_{n}^{\downarrow}$ such that
(7) $\hat{\rho}_{n}(x)=\sum_{i=1}^{n}a_{i}(-x_{i:n}),\qquad x\in\mathbb{R}^{n}.$
Our goal in the following sections is to provide a unified hyperfinite perspective on these results and to establish new consistency and asymptotic theorems.
3. Non-standard analysis: a self-contained introduction
This section provides the essential non-standard machinery required for our applications. We aim for a self-contained exposition accessible to readers without prior exposure to NSA, whilst maintaining sufficient rigour for our proofs. For comprehensive treatments, see Keisler [13], Albeverio et al. [3], and Fajardo–Keisler [10].
3.1. Standing conventions and the Loeb model
Convention 3.1 (Saturation assumption).
Throughout this paper, we work in a countably saturated non-standard enlargement: every countable collection of internal sets with the finite intersection property has non-empty intersection. This level of saturation suffices for all constructions in this paper, including overspill/underspill arguments and the existence of Loeb measures. Countable saturation can be arranged in standard constructions, e.g., via an ultrapower by a non-principal ultrafilter on $\mathbb{N}$ in a countable superstructure setting; see [3, Ch. 4].
Notation 3.2 (Standing Loeb model).
Throughout Sections 310 we work inside a fixed countably saturated non-standard universe. When we invoke hyperfinite probability, we use the following canonical model.
We fix an infinite $N\in{}^{*}\mathbb{N}$, chosen once and for all by the mechanism described in Lemma 3.3 below. The hyperfinite set $I_{N}=\{1,\dots,N\}$ carries the internal counting measure $\mu_{N}(A)=|A|/N$ on the algebra $\mathscr{I}_{N}$ of internal subsets. Let $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ denote the associated Loeb probability space.
When we state results about i.i.d. samples $(X_{i})$ from a random variable $X$ on an ambient probability space $(\Omega,\mathscr{G},\mathsf{P})$, we implicitly pass to the non-standard extension $({}^{*}\Omega,{}^{*}\mathscr{G},{}^{*}\mathsf{P})$ and its Loeb completion $({}^{*}\Omega,L({}^{*}\mathscr{G}),L({}^{*}\mathsf{P}))$; all “almost sure” statements in NSA arguments are understood with respect to $L({}^{*}\mathsf{P})$ unless explicitly stated otherwise.
Star-suppression convention. When a standard function $f$ is evaluated at a hyperreal argument, we mean, by convention, its non-standard extension $^{*}f$; e.g. $q(\alpha_{k})$ abbreviates $^{*}q(\alpha_{k})$.
Lemma 3.3 (Simultaneous validity of hyperfinite properties).
Let $(P_{j})_{j\in\mathbb{N}}$ be a countable family of properties, where each $P_{j}(n)$ is a first-order statement about sample size $n$ expressible with rational parameters (e.g., $^{*}\mathsf{P}(|\cdots|>\varepsilon)<\delta$” for rational $\varepsilon,\delta>0$). By the internal definition principle, each set
$S_{j}:=\{n\in{}^{*}\mathbb{N}:P_{j}(n)\text{ holds}\}$
is internal. Assume each $P_{j}$ holds eventually for standard $n$, i.e., $S_{j}$ contains all sufficiently large standard naturals.
Then there exists an infinite $N\in{}^{*}\mathbb{N}$ such that $P_{j}(N)$ holds for all $j\in\mathbb{N}$.
Proof.
For each standard $m\in\mathbb{N}$, define the internal set $B_{m}:=\{n\in{}^{*}\mathbb{N}:n>m\}$. Consider the countable family of internal sets $\{S_{j}\}_{j\in\mathbb{N}}\cup\{B_{m}\}_{m\in\mathbb{N}}$.
This family has the finite intersection property: for any finite $J\subset\mathbb{N}$ and finite $M\subset\mathbb{N}$, the intersection $\bigcap_{j\in J}S_{j}\cap\bigcap_{m\in M}B_{m}$ contains all standard $n>\max(\max_{j\in J}n_{0}^{(j)},\max M)$, hence is non-empty.
By countable saturation, $\bigcap_{j\in\mathbb{N}}S_{j}\cap\bigcap_{m\in\mathbb{N}}B_{m}\neq\emptyset$. Any element $N$ of this intersection satisfies $P_{j}(N)$ for all $j$ (since $N\in S_{j}$) and $N>m$ for all standard $m$ (since $N\in B_{m}$), hence $N$ is infinite. ∎
Remark 3.4 (Properties used in this paper).
The properties $P_{j}$ we invoke include: the hyperfinite strong law of large numbers (Theorem 3.18) for countably many integrands arising from truncations, the hyperfinite Glivenko–Cantelli theorem (Theorem 3.20), and the hyperfinite CLT (Theorem 10.1). Each is expressed in the first-order form $\forall\varepsilon\in\mathbb{Q}_{>0}\,\exists n_{0}\,\forall n\geqslant n_{0}$: [bound holds]”, which transfers to give internal sets $S_{j}$. Lemma 3.3 then guarantees a single infinite $N$ for which all hold simultaneously.
Proposition 3.5 (Atomless separable spaces and Loeb spaces).
Every atomless separable (i.e., countably generated mod null sets) standard probability space $(\Omega,\mathscr{G},\mathsf{P})$ is measure-algebra isomorphic to a hyperfinite Loeb space $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ for some infinite $N\in{}^{*}\mathbb{N}$. Precisely, there is a Boolean algebra isomorphism
$\Phi:\mathscr{G}/\mathcal{N}_{\mathsf{P}}\xrightarrow{\;\cong\;}\mathscr{I}_{N}^{L}/\mathcal{N}_{L(\mu_{N})}$
between the measure algebras (quotients by null sets) such that $\mathsf{P}(A)=L(\mu_{N})(\Phi([A]))$ for all $A\in\mathscr{G}$. This isomorphism induces isometric lattice isomorphisms $L^{p}(\Omega,\mathsf{P})\cong L^{p}(I_{N},L(\mu_{N}))$ for all $p\in[1,\infty]$.
Proof (reference).
This is Keisler’s representation theorem; see [10, Thm. 10.3.1] or [3, Ch. 5]. The key fact is that the hyperfinite Loeb space $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ with infinite $N$ is atomless and has the same measure algebra (mod null sets) as the Lebesgue unit interval. Every atomless separable probability space shares this property by Maharam’s theorem. ∎
For distributional/statistical arguments, when the underlying standard probability space is atomless and separable (which includes all standard Borel probability spaces), we may replace it by an isomorphic Loeb space via Proposition 3.5. This convention avoids repeated measure-theoretic bookkeeping and allows us to treat hyperfinite sums as Loeb integrals via standard part.
3.2. The hyperreal numbers
The starting point of non-standard analysis is the construction of an ordered field extension $^{*}\mathbb{R}$ of $\mathbb{R}$, called the hyperreals. There are several approaches to this construction (ultraproducts, superstructures, axiomatic); we adopt the axiomatic viewpoint, which suffices for applications.
Definition 3.6 (Hyperreal numbers).
The hyperreal numbers $^{*}\mathbb{R}$ form an ordered field extension of $\mathbb{R}$ satisfying the following properties:
  1. (H1)
    $\mathbb{R}\subset{}^{*}\mathbb{R}$ is a proper ordered subfield.
  2. (H2)
    (Transfer principle) Every first-order statement about $\mathbb{R}$ that is true remains true when interpreted in $^{*}\mathbb{R}$.
  3. (H3)
    There exist infinitesimals $\varepsilon\in{}^{*}\mathbb{R}$ with $|\varepsilon|<1/n$ for all $n\in\mathbb{N}$.
  4. (H4)
    There exist infinite hyperreals $H\in{}^{*}\mathbb{R}$ with $|H|>n$ for all $n\in\mathbb{N}$.
(In standard constructions, (H3) and (H4) follow from (H1) and (H2) together with properness.)
The transfer principle is the key tool for moving between standard and non-standard contexts. It asserts that the hyperreals satisfy the same first-order properties as the reals. For instance, the statement “for all $x,y\in\mathbb{R}$ with $x<y$, there exists $z\in\mathbb{R}$ with x<z<yx<z<y” transfers to “for all $x,y\in{}^{*}\mathbb{R}$ with $x<y$, there exists $z\in{}^{*}\mathbb{R}$ with x<z<yx<z<y.”
Definition 3.7 (Finite, infinitesimal, infinite).
A hyperreal $x\in{}^{*}\mathbb{R}$ is:
  • finite if $|x|\leqslant n$ for some $n\in\mathbb{N}$;
  • infinitesimal if $|x|<1/n$ for all $n\in\mathbb{N}$;
  • infinite if $|x|>n$ for all $n\in\mathbb{N}$.
Two hyperreals $x,y$ are infinitely close, written $x\approx y$, if xyx-y is infinitesimal.
The following proposition provides the crucial connection between finite hyperreals and real numbers.
Proposition 3.8 (Standard part).
Every finite hyperreal $x\in{}^{*}\mathbb{R}$ is infinitely close to a unique real number, denoted $\operatorname{st}(x)\in\mathbb{R}$ and called its standard part. The map $\operatorname{st}:\{x\in{}^{*}\mathbb{R}:x\text{ is finite}\}\to\mathbb{R}$ satisfies:
  1. (i)
    $\operatorname{st}(x+y)=\operatorname{st}(x)+\operatorname{st}(y)$ for finite $x,y$;
  2. (ii)
    $\operatorname{st}(xy)=\operatorname{st}(x)\operatorname{st}(y)$ for finite $x,y$;
  3. (iii)
    $\operatorname{st}(x)\leqslant\operatorname{st}(y)$ if $x\leqslant y$ for finite $x,y$;
  4. (iv)
    $\operatorname{st}(r)=r$ for all $r\in\mathbb{R}$.
Proof.
For any finite $x$, the set $\{r\in\mathbb{R}:r\leqslant x\}$ is non-empty and bounded above, hence has a supremum $s:=\sup\{r\in\mathbb{R}:r\leqslant x\}$. We claim $x\approx s$. If xs>1/nx-s>1/n for some $n\in\mathbb{N}$, then $s+1/(2n)<x$ and $s+1/(2n)\in\mathbb{R}$, contradicting the definition of $s$. Similarly, if sx>1/ns-x>1/n, then s1/(2n)>xs-1/(2n)>x implies s1/(2n)s-1/(2n) is an upper bound for $\{r\in\mathbb{R}:r\leqslant x\}$, contradicting the supremum. Thus |xs|<1/n|x-s|<1/n for all $n$, i.e., $x\approx s$.
Uniqueness: if $x\approx s$ and $x\approx s^{\prime}$ with $s,s^{\prime}\in\mathbb{R}$, then $s-s^{\prime}$ is infinitesimal and real, hence zero.
The algebraic properties follow from the corresponding properties of $\approx$ and the field operations. ∎
3.3. The hypernaturals and hyperfinite sets
The natural numbers $\mathbb{N}$ extend to the hypernaturals $^{*}\mathbb{N}\subset{}^{*}\mathbb{R}$. By the transfer principle, $^{*}\mathbb{N}$ satisfies all first-order properties of $\mathbb{N}$. However, $^{*}\mathbb{N}$ properly contains $\mathbb{N}$: there exist infinite hypernaturals $N\in{}^{*}\mathbb{N}\setminus\mathbb{N}$ satisfying $N>n$ for all $n\in\mathbb{N}$.
Definition 3.9 (Hyperfinite set).
For $N\in{}^{*}\mathbb{N}$, the set
$I_{N}:=\{1,2,\ldots,N\}=\{k\in{}^{*}\mathbb{N}:1\leqslant k\leqslant N\}$
is called a hyperfinite set. It is internal in the sense that it arises from the non-standard extension.
The crucial property of hyperfinite sets is that they behave like finite sets from the internal perspective. In particular:
Proposition 3.10 (Internal finite operations).
Let $N\in{}^{*}\mathbb{N}$ (possibly infinite) and let $f:I_{N}\to{}^{*}\mathbb{R}$ be an internal function. Then:
  1. (i)
    The hyperfinite sum $\sum_{k=1}^{N}f(k)\in{}^{*}\mathbb{R}$ is well-defined.
  2. (ii)
    The hyperfinite product $\prod_{k=1}^{N}f(k)\in{}^{*}\mathbb{R}$ is well-defined.
  3. (iii)
    The maximum $\max_{k\in I_{N}}f(k)$ and minimum $\min_{k\in I_{N}}f(k)$ are well-defined.
  4. (iv)
    Sorting: if $f$ takes values in $^{*}\mathbb{R}$, then there exists an internal permutation $\sigma:I_{N}\to I_{N}$ such that $f(\sigma(1))\leqslant f(\sigma(2))\leqslant\cdots\leqslant f(\sigma(N))$.
These properties follow from the transfer principle applied to the corresponding statements about finite sets. In particular, for (iv) we transfer the statement “for every finite set $S$ and function $g:S\to\mathbb{R}$ there exists a permutation $\sigma$ sorting gg”; this yields an internal permutation $\sigma$, which is essential for later applications to order statistics.
3.4. Internal and external sets
A fundamental distinction in NSA is between internal and external sets. Internal sets arise from the non-standard extension and satisfy the transfer principle; external sets do not.
Example 3.11.
The set $I_{N}=\{1,2,\dots,N\}$ for $N\in{}^{*}\mathbb{N}$ is internal. The set $\mathbb{N}$ itself, viewed as a subset of $^{*}\mathbb{N}$, is external when $^{*}\mathbb{N}\neq\mathbb{N}$.
A convenient way to see this without appealing to second-order quantification is the following. In the standard universe, for every $m\in\mathbb{N}$ every non-empty subset of $\{1,\dots,m\}$ has a maximum. This is a first-order statement in the superstructure. By transfer, for every $M\in{}^{*}\mathbb{N}$ every non-empty internal subset of $I_{M}$ has a maximum.
Now pick an infinite $H\in{}^{*}\mathbb{N}$. Then $\mathbb{N}\subseteq I_{H}$. If $\mathbb{N}$ were internal, it would be an internal non-empty subset of $I_{H}$ and hence would have a maximum, contradicting the fact that $\mathbb{N}$ has no largest element. Therefore $\mathbb{N}$ is external.
Proposition 3.12 (Internal definition principle).
If $P(x)$ is a first-order property and $A$ is an internal set, then $\{x\in A:P(x)\}$ is internal.
3.5. Loeb measure: from hyperfinite to standard probability
The Loeb construction [15] is the key tool for producing genuine probability spaces from internal hyperfinite structures. We describe it in the context most relevant to our applications.
Fix an infinite $N\in{}^{*}\mathbb{N}$ and consider the hyperfinite set $I_{N}=\{1,\ldots,N\}$. Let $\mathscr{I}_{N}$ denote the algebra of all internal subsets of $I_{N}$. Define the internal counting probability measure $\mu_{N}:\mathscr{I}_{N}\to{}^{*}[0,1]$ by
(8) $\mu_{N}(A):=\frac{|A|}{N},\qquad A\in\mathscr{I}_{N},$
where $|A|\in{}^{*}\mathbb{N}$ is the internal cardinality of $A$.
The triple $(I_{N},\mathscr{I}_{N},\mu_{N})$ is an internal finitely additive probability space. The Loeb construction promotes this to a genuine $\sigma$-additive probability space.
Theorem 3.13 (Loeb measure construction).
Define the Loeb premeasure $\mu_{N}^{0}:\mathscr{I}_{N}\to[0,1]$ by $\mu_{N}^{0}(A):=\operatorname{st}(\mu_{N}(A))$ for $A\in\mathscr{I}_{N}$. Define an outer measure $L^{*}(\mu_{N})$ on all subsets $B\subseteq I_{N}$ by
$L^{*}(\mu_{N})(B):=\inf\{\mu_{N}^{0}(A):\ A\in\mathscr{I}_{N},\ B\subseteq A\}.$
Let $\mathscr{I}_{N}^{L}$ be the $\sigma$-algebra of $L^{*}(\mu_{N})$-measurable sets in the sense of Carathéodory, and define $L(\mu_{N}):=L^{*}(\mu_{N})\!\upharpoonright_{\mathscr{I}_{N}^{L}}$. Then:
  1. (i)
    $\mu_{N}^{0}$ is a finitely additive probability measure on the internal algebra $\mathscr{I}_{N}$.
  2. (ii)
    $\mathscr{I}_{N}\subseteq\mathscr{I}_{N}^{L}$, and $L(\mu_{N})$ is a complete $\sigma$-additive probability measure on $\mathscr{I}_{N}^{L}$ extending $\mu_{N}^{0}$.
  3. (iii)
    In particular, $\sigma(\mathscr{I}_{N})\subseteq\mathscr{I}_{N}^{L}$, and the restriction of $L(\mu_{N})$ to $\sigma(\mathscr{I}_{N})$ is a $\sigma$-additive extension of $\mu_{N}^{0}$.
Proof (reference).
The key step is verifying that the Carathéodory construction yields a complete $\sigma$-additive measure. The countable subadditivity of $\mu_{N}^{0}$ on $\mathscr{I}_{N}$ (which follows from the internal finite additivity and saturation) is the crucial ingredient. For the full argument, see Loeb [15] or [3, Ch. 5] and [10, Chs. 4–5]. ∎
Remark 3.14.
The Loeb space is rich enough to support continuous distributions, despite being built from a discrete internal structure. For instance, if $N$ is infinite, the Loeb space $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ is atomless.
3.6. Hyperfinite integration and the Riemann sum approximation
The Loeb measure allows us to integrate functions on hyperfinite spaces. For our purposes, the following connection between hyperfinite sums and standard integrals is fundamental.
Proposition 3.15 (Hyperfinite Riemann approximation).
Let $f:[0,1]\to\mathbb{R}$ be bounded and Riemann integrable, and let $^{*}f:{}^{*}[0,1]\to{}^{*}\mathbb{R}$ be its non-standard extension. For any infinite $N\in{}^{*}\mathbb{N}$,
(9) $\int_{0}^{1}f(t)\,dt=\operatorname{st}\left(\frac{1}{N}\sum_{k=1}^{N}{}^{*}f\left(\frac{k}{N}\right)\right).$
Proof.
By Riemann integrability, for each $\varepsilon>0$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
$\left|\int_{0}^{1}f(t)\,dt-\frac{1}{n}\sum_{k=1}^{n}f\left(\frac{k}{n}\right)\right|<\varepsilon.$
By transfer, this statement holds for all $n\in{}^{*}\mathbb{N}$ with $n\geqslant n_{0}$. In particular, it holds for the infinite $N$, giving
$\left|\int_{0}^{1}f(t)\,dt-\frac{1}{N}\sum_{k=1}^{N}{}^{*}f\left(\frac{k}{N}\right)\right|<\varepsilon.$
Since $\varepsilon>0$ was arbitrary, the hyperfinite sum differs from the integral by an infinitesimal, and taking standard parts yields (9). ∎
Proposition 3.16 (Loeb lifting for $L^{1}$-functions on the hyperfinite grid).
Let $f\in L^{1}([0,1],\lambda)$, where $\lambda$ is Lebesgue measure. Fix an infinite $N\in{}^{*}\mathbb{N}$ and write $I_{N}=\{1,\dots,N\}$. Then there exists an internal function $F:I_{N}\to{}^{*}\mathbb{R}$ such that:
  1. (i)
    $f(\operatorname{st}(k/N))=\operatorname{st}(F(k))$ for $L(\mu_{N})$-almost all $k\in I_{N}$;
  2. (ii)
    $F$ is $S$-integrable (equivalently, Loeb integrable) and
    $\int_{0}^{1}f(t)\,dt=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}F(k)\Big).$
If, in addition, $f$ is bounded and Riemann integrable, one may choose $F(k)={}^{*}f(k/N)$.
Proof (reference).
This is a standard lifting result in Loeb integration theory: every $L^{1}$ function admits an internal $S$-integrable lifting on the hyperfinite grid, and the Loeb integral equals the standard part of the internal counting integral. See, for example, [10, Chs. 4–5] (Loeb liftings and $S$-integrability) or [3, Ch. 5]. The last sentence follows from Proposition 3.15. ∎
Remark 3.17 (A concrete $S$-integrability criterion).
An internal function $F:I_{N}\to{}^{*}\mathbb{R}$ is called $S$-integrable if $\frac{1}{N}\sum_{k=1}^{N}|F(k)|$ is finite and, for every internal $A\subseteq I_{N}$ with $\mu_{N}(A)\approx 0$, one has $\frac{1}{N}\sum_{k\in A}|F(k)|\approx 0$. In that case $\operatorname{st}(F)$ is Loeb integrable and its Loeb integral equals the standard part of the internal counting integral.
3.7. The hyperfinite strong law and Glivenko–Cantelli theorem
For statistical applications, we need the non-standard versions of classical limit theorems.
Theorem 3.18 (Hyperfinite strong law of large numbers).
Let $(X_{i})_{i\in\mathbb{N}}$ be i.i.d. with $\mathsf{E}[|X_{1}|]<\infty$ on $(\Omega,\mathscr{G},\mathsf{P})$. Let $N\in{}^{*}\mathbb{N}$ be infinite and consider the hyperfinite extension $(X_{1},\dots,X_{N})$ as random variables on $({}^{*}\Omega,L({}^{*}\mathscr{G}),L({}^{*}\mathsf{P}))$. Then
$\frac{1}{N}\sum_{k=1}^{N}X_{k}\approx\mathsf{E}[X_{1}]\qquad\text{holds $L({}^{*}\mathsf{P})$-almost surely.}$
Proof.
By the classical strong law, $\frac{1}{n}\sum_{k=1}^{n}X_{k}\to\mathsf{E}[X_{1}]$ almost surely as $n\to\infty$. For each rational $\varepsilon>0$, define
$A_{\varepsilon}:=\left\{\omega\in\Omega:\exists n_{0}\in\mathbb{N}\text{ such that }\forall n\geqslant n_{0},\left|\frac{1}{n}\sum_{k=1}^{n}X_{k}(\omega)-\mathsf{E}[X_{1}]\right|<\varepsilon\right\}.$
This set is $\mathscr{G}$-measurable (it is a countable union/intersection of measurable events), and the strong law gives $\mathsf{P}(A_{\varepsilon})=1$.
The non-standard extension $^{*}A_{\varepsilon}\subseteq{}^{*}\Omega$ is internal (it is the star-extension of a standard set), hence $^{*}A_{\varepsilon}\in{}^{*}\mathscr{G}\subseteq L({}^{*}\mathscr{G})$. Moreover,
$L({}^{*}\mathsf{P})({}^{*}A_{\varepsilon})=\operatorname{st}({}^{*}\mathsf{P}({}^{*}A_{\varepsilon}))=\operatorname{st}(\mathsf{P}(A_{\varepsilon}))=1,$
using the fact that $^{*}\mathsf{P}({}^{*}A_{\varepsilon})=\mathsf{P}(A_{\varepsilon})$ by transfer of the probability.
For $\omega\in{}^{*}A_{\varepsilon}$, the internal statement $\exists n_{0}\in{}^{*}\mathbb{N}$ such that $\forall n\in{}^{*}\mathbb{N}$ with $n\geqslant n_{0}$, $|\frac{1}{n}\sum_{k=1}^{n}X_{k}(\omega)-\mathsf{E}[X_{1}]|<\varepsilon$” holds (by transfer of the defining property of $A_{\varepsilon}$). In particular, it holds for the infinite $N$, giving $|\frac{1}{N}\sum_{k=1}^{N}X_{k}(\omega)-\mathsf{E}[X_{1}]|<\varepsilon$.
Since $L({}^{*}\mathsf{P})({}^{*}A_{\varepsilon})=1$ for each rational $\varepsilon>0$, intersecting over $\varepsilon\in\mathbb{Q}_{>0}$ (a countable intersection of Loeb-measure-one sets) yields the claim $L({}^{*}\mathsf{P})$-almost surely. ∎
Remark 3.19 (Two-level almost sure statements).
In hyperfinite sampling statements we use two measures simultaneously: $L({}^{*}\mathsf{P})$ governs the randomness of the sample path $\omega\in{}^{*}\Omega$, while $L(\mu_{N})$ governs the fraction of indices $k\in I_{N}$ for a fixed sample path. Thus “for $L(\mu_{N})$-almost all kk” should be read as: for $L({}^{*}\mathsf{P})$-almost all sample outcomes, the exceptional set of indices has Loeb counting measure zero.
Theorem 3.20 (Hyperfinite Glivenko–Cantelli / quantile shadow).
Let $(X_{i})_{i\in\mathbb{N}}$ be i.i.d. with distribution function $F$ and lower quantile function $q(\alpha)=\inf\{x:\,F(x)\geqslant\alpha\}$. Assume $\mathsf{E}[|X_{1}|]<\infty$, equivalently $\int_{0}^{1}|q(\alpha)|\,d\alpha<\infty$.
Fix an infinite $N\in{}^{*}\mathbb{N}$ and consider the hyperfinite sample $(X_{1},\dots,X_{N})$ with order statistics $X_{1:N}\leqslant\cdots\leqslant X_{N:N}$. Let $\alpha_{k}:=k/N$.
Then $L({}^{*}\mathsf{P})$-almost surely, the set of indices
$G:=\{k\in I_{N}:\ X_{k:N}\approx q(\alpha_{k})\}$
has Loeb counting measure $L(\mu_{N})(G)=1$.
Moreover, for every bounded Riemann integrable $g:[0,1]\to\mathbb{R}$ with $\int_{0}^{1}|g(\alpha)q(\alpha)|\,d\alpha<\infty$, we have
(10) $\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}g(\alpha_{k})X_{k:N}\Big)=\int_{0}^{1}g(\alpha)q(\alpha)\,d\alpha.$
Proof.
Internal infinitesimal bound via convergence in probability. The classical Glivenko–Cantelli theorem states that $\sup_{x}|F_{n}(x)-F(x)|\to 0$ almost surely. The following weaker consequence suffices: for every $\eta>0$,
$\mathsf{P}\Big(\sup_{x}|F_{n}(x)-F(x)|>\eta\Big)\to 0\quad\text{as }n\to\infty.$
This is a first-order statement: for every standard $\eta>0$ and $\delta>0$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$, $\mathsf{P}(\sup_{x}|F_{n}-F|>\eta)<\delta$.
By transfer, for our fixed infinite $N$ and any standard $\eta,\delta>0$,
$^{*}\mathsf{P}\Big(\sup_{x}|F_{N}(x)-F(x)|>\eta\Big)<\delta.$
Taking Loeb measures (standard parts), $L({}^{*}\mathsf{P})(\sup_{x}|F_{N}-F|>\eta)\leqslant\delta$. Since $\delta>0$ was arbitrary, $L({}^{*}\mathsf{P})(\sup_{x}|F_{N}-F|>\eta)=0$.
Intersecting over a countable sequence $\eta=1/m$ for $m\in\mathbb{N}$, we obtain
$L({}^{*}\mathsf{P})\Big(\sup_{x}|F_{N}(x)-F(x)|\text{ is infinitesimal}\Big)=1.$
Fix a sample path in this Loeb-probability-one event, and let $\varepsilon_{0}:=\sup_{x}|F_{N}(x)-F(x)|$. Then $\varepsilon_{0}\approx 0$ is a positive infinitesimal. Moreover, $\varepsilon_{0}$ is internal (it is the internal supremum of an internal function), so we may use it in transferred statements.
The standard quantile bracketing implication: for any $\alpha\in(0,1)$ and $\varepsilon\in(0,\alpha\wedge(1-\alpha))$,
(11) $\sup_{x}|F_{n}(x)-F(x)|\leqslant\varepsilon\quad\Longrightarrow\quad q(\alpha-\varepsilon)\leqslant q_{n}(\alpha)\leqslant q(\alpha+\varepsilon).$
This can be expressed using only bounded quantification over $\mathbb{R}$ and the defining formula for the lower quantile, so it is a first-order statement in $\alpha$, $\varepsilon$, and $F_{n}$. By transfer, (11) holds for all hyperreal $\alpha\in{}^{*}(0,1)$ and hyperreal $\varepsilon\in{}^{*}(0,\alpha\wedge(1-\alpha))$ with $\|F_{N}-F\|_{\infty}\leqslant\varepsilon$.
Since $\|F_{N}-F\|_{\infty}\leqslant\varepsilon_{0}$ and $\varepsilon_{0}$ is infinitesimal, for any hyperreal $\alpha\in{}^{*}(0,1)$ with $\alpha>\varepsilon_{0}$ and $\alpha<1-\varepsilon_{0}$,
$q(\alpha-\varepsilon_{0})\leqslant q_{N}(\alpha)\leqslant q(\alpha+\varepsilon_{0}).$
In particular, $q_{N}(\alpha_{k})\leqslant q(\alpha_{k}+\varepsilon_{0})$ and $q_{N}(\alpha_{k})\geqslant q(\alpha_{k}-\varepsilon_{0})$.
Halos of discontinuities and boundary have Loeb measure zero. Since $q$ is monotone, its set of discontinuity points $\mathrm{Disc}(q)\subseteq(0,1)$ is at most countable. For each standard $\alpha_{0}\in\mathrm{Disc}(q)\cup\{0,1\}$, define the halo
$H_{\alpha_{0}}:=\{k\in I_{N}:\ \alpha_{k}\approx\alpha_{0}\}=\{k\in I_{N}:\ |\alpha_{k}-\alpha_{0}|\ \text{is infinitesimal}\}.$
We claim $L(\mu_{N})(H_{\alpha_{0}})=0$. For any standard $\delta>0$,
$H_{\alpha_{0}}\subseteq\{k\in I_{N}:\ |\alpha_{k}-\alpha_{0}|<\delta\},$
and the right-hand side is internal with
$\mu_{N}(\{k:|\alpha_{k}-\alpha_{0}|<\delta\})\leqslant 2\delta.$
Taking standard parts, $L(\mu_{N})(H_{\alpha_{0}})\leqslant 2\delta$. Since $\delta>0$ was arbitrary, $L(\mu_{N})(H_{\alpha_{0}})=0$. (Note: $H_{\alpha_{0}}$ has outer Loeb measure zero, hence is Loeb measurable.)
Define
$D:=\bigcup_{\alpha_{0}\in\mathrm{Disc}(q)\cup\{0,1\}}H_{\alpha_{0}}.$
Since $\mathrm{Disc}(q)\cup\{0,1\}$ is countable and each $H_{\alpha_{0}}$ has Loeb measure 0, and $L(\mu_{N})$ is countably additive, $L(\mu_{N})(D)=0$.
The quantile shadow on $G=I_{N}\setminus D$. Take $k\notin D$, so $\alpha_{k}$ is not in the halo of any discontinuity of $q$ nor in the halo of $\{0,1\}$. This means $t:=\operatorname{st}(\alpha_{k})\in(0,1)\setminus\mathrm{Disc}(q)$, hence $q$ is continuous at $t$.
Since $q$ is continuous at $t$ and $\alpha_{k}\approx t$, both $\alpha_{k}-\varepsilon_{0}\approx t$ and $\alpha_{k}+\varepsilon_{0}\approx t$ (because $\varepsilon_{0}$ is infinitesimal). Therefore
$q(\alpha_{k}-\varepsilon_{0})\approx q(t)\approx q(\alpha_{k}+\varepsilon_{0}).$
The transferred bracketing gives
$q(\alpha_{k}-\varepsilon_{0})\leqslant q_{N}(\alpha_{k})=X_{k:N}\leqslant q(\alpha_{k}+\varepsilon_{0}),$
hence $X_{k:N}\approx q(t)=q(\operatorname{st}(\alpha_{k}))\approx q(\alpha_{k})$ (using continuity again). Thus $k\in G$, and we conclude $G\supseteq I_{N}\setminus D$, hence $L(\mu_{N})(G)=1$.
Convention for the integral identity. We interpret integrals involving internal functions (such as $q_{N}$ and $g_{N}$) as Loeb integrals on the hyperfinite grid $\{\alpha_{k}:k\in I_{N}\}$, so that
$\int_{0}^{1}H(\alpha)\,d\alpha=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}H(\alpha_{k})\Big)$
whenever $H$ is $S$-integrable on the grid (cf. Remark 3.17 and Lemma 4.1).
The integral identity via truncation. Define the step function $g_{N}(\alpha):=g(\alpha_{k})$ on ((k1)/N,k/N]((k-1)/N,k/N]. Then
$\frac{1}{N}\sum_{k=1}^{N}g(\alpha_{k})X_{k:N}=\int_{0}^{1}g_{N}(\alpha)\,q_{N}(\alpha)\,d\alpha$
because $q_{N}(\alpha)=X_{k:N}$ on ((k1)/N,k/N]((k-1)/N,k/N].
We prove $\int_{0}^{1}|q_{N}-q|\,d\alpha\approx 0$ directly by truncation (not as a consequence of the quantile shadow, which only gives pointwise information on a full-Loeb set). For $M>0$, define $q^{(M)}:=\max(\min(q,M),-M)$ and $q_{N}^{(M)}:=\max(\min(q_{N},M),-M)$. Then
$\int_{0}^{1}|q_{N}-q|\leqslant\int_{0}^{1}|q_{N}^{(M)}-q^{(M)}|+\int_{0}^{1}|q_{N}-q_{N}^{(M)}|+\int_{0}^{1}|q-q^{(M)}|.$
For the first term: $|q_{N}^{(M)}-q^{(M)}|\leqslant 2M$ pointwise, and on the full-Loeb set $G$, $q_{N}^{(M)}(\alpha_{k})\approx q^{(M)}(\alpha_{k})$. By the Loeb dominated convergence theorem (see [3, Thm. 4.3.6]: for internal functions dominated by an $S$-integrable bound, pointwise near-equality on a Loeb-measure-one set implies the integrals are infinitely close), $\int_{0}^{1}|q_{N}^{(M)}-q^{(M)}|\approx 0$ for each fixed standard $M$.
For the second term: $|q_{N}-q_{N}^{(M)}|=|q_{N}|\mathbf{1}_{|q_{N}|>M}$. Since $\mathsf{E}|X_{1}|<\infty$, the hyperfinite SLLN gives $\frac{1}{N}\sum_{k=1}^{N}|X_{k}|\approx\mathsf{E}|X_{1}|<\infty$. By the standard integration identity for order statistics, $\int_{0}^{1}|q_{N}|\,d\alpha=\frac{1}{N}\sum_{k=1}^{N}|X_{k:N}|$, which has finite standard part. Thus $\int_{0}^{1}|q_{N}-q_{N}^{(M)}|\to 0$ as $M\to\infty$ (in the sense that for any standard $\delta>0$, there exists standard $M$ such that this integral $\delta$).
For the third term: $\int_{0}^{1}|q-q^{(M)}|\to 0$ as $M\to\infty$ by $\mathsf{E}|X_{1}|<\infty$.
Given $\delta>0$, choose $M$ large enough that the second and third terms are each $\delta/3$. Then choose the infinitesimal such that the first term is infinitesimal. Taking standard parts, $\operatorname{st}(\int_{0}^{1}|q_{N}-q|)\leqslant 2\delta/3<\delta$. Since $\delta$ was arbitrary, $\int_{0}^{1}|q_{N}-q|\approx 0$.
Now we can complete the proof of (10):
$\int_{0}^{1}g_{N}q_{N}-\int_{0}^{1}gq=\int_{0}^{1}g_{N}(q_{N}-q)+\int_{0}^{1}(g_{N}-g)q.$
The first term is bounded by $\|g\|_{\infty}\int_{0}^{1}|q_{N}-q|\approx 0$. For the second term, fix $M$ and split $q=q^{(M)}+(q-q^{(M)})$:
$\Big|\int_{0}^{1}(g_{N}-g)q\Big|\leqslant\|q^{(M)}\|_{\infty}\int_{0}^{1}|g_{N}-g|+2\|g\|_{\infty}\int_{0}^{1}|q-q^{(M)}|.$
Since $g$ is Riemann integrable, $\int_{0}^{1}|g_{N}-g|\approx 0$ (by the hyperfinite Riemann sum approximation). Let $M\to\infty$ and use $\int_{0}^{1}|q-q^{(M)}|\to 0$. Taking standard parts yields (10). ∎
3.8. Overspill and underspill
Two technical principles are frequently used in NSA arguments. We assume throughout that our non-standard universe is countably saturated: every countable collection of internal sets with the finite intersection property has non-empty intersection.
Proposition 3.21 (Overspill).
Let $A\subseteq{}^{*}\mathbb{N}$ be internal. If $A$ contains all standard natural numbers, then $A$ contains some infinite hypernatural.
Proof.
For each standard $m\in\mathbb{N}$, define the internal set
$A_{m}:=\{n\in{}^{*}\mathbb{N}:\ n\in A\ \wedge\ n>m\}.$
Each $A_{m}$ is internal by the internal definition principle (intersection of $A$ with the internal set $\{n\in{}^{*}\mathbb{N}:n>m\}$). Each $A_{m}$ is non-empty since $m+1\in A$ (as $A$ contains all standard naturals) and $m+1>m$.
The collection $\{A_{m}:m\in\mathbb{N}\}$ is countable and has the finite intersection property: for any $m_{1},\dots,m_{k}\in\mathbb{N}$, we have $A_{m_{1}}\cap\cdots\cap A_{m_{k}}\supseteq A_{\max(m_{1},\dots,m_{k})}\neq\emptyset$.
By countable saturation, $\bigcap_{m\in\mathbb{N}}A_{m}\neq\emptyset$. Pick $H$ in this intersection. Then $H\in A$ and $H>m$ for every standard $m\in\mathbb{N}$, hence $H$ is infinite. ∎
Proposition 3.22 (Underspill / underflow).
Let $A\subseteq{}^{*}\mathbb{N}$ be internal. If $A$ contains every infinite hypernatural, then $A$ contains some standard natural number.
Equivalently: if $A\cap\mathbb{N}=\emptyset$, then there exists an infinite $H\in{}^{*}\mathbb{N}$ such that $A\subseteq\{H+1,H+2,\dots\}$.
Proof.
Assume first that $A$ contains every infinite hypernatural and set $B:={}^{*}\mathbb{N}\setminus A$. Then $B$ is internal. If $B=\emptyset$, then $A={}^{*}\mathbb{N}$ and in particular $A$ contains every standard natural, so the conclusion holds. Hence we may assume $B\neq\emptyset$.
By assumption, $B$ contains no infinite hypernaturals, hence every element of $B$ is finite and therefore standard; thus $B\subseteq\mathbb{N}$.
Fix an infinite $H\in{}^{*}\mathbb{N}$. Then $B\subseteq I_{H}=\{1,\dots,H\}$. Since $B$ is a non-empty internal subset of the hyperfinite set $I_{H}$, transfer of the finite-maximum principle implies that $B$ has a maximum element $m=\max B$. This $m$ cannot be infinite (because $B$ contains no infinite hypernaturals), hence $m\in\mathbb{N}$. Therefore $B\subseteq\{1,\dots,m\}$, so $B$ is finite. Consequently $A$ contains all standard naturals >m>m, and in particular $A$ contains some standard natural number.
For the equivalent formulation, assume $A\cap\mathbb{N}=\emptyset$, so $\mathbb{N}\subseteq{}^{*}\mathbb{N}\setminus A$. Consider the internal set
$C:=\{m\in{}^{*}\mathbb{N}:\ \{1,\dots,m\}\subseteq{}^{*}\mathbb{N}\setminus A\}.$
Then $C$ contains every standard $m\in\mathbb{N}$, hence by overspill $C$ contains an infinite $H$. Thus $\{1,\dots,H\}\subseteq{}^{*}\mathbb{N}\setminus A$, which is equivalent to $A\subseteq\{H+1,H+2,\dots\}$. ∎
4. Hyperfinite representation of coherent risk measures
We now develop the hyperfinite representation of coherent risk measures, which forms the foundation for our unified treatment of CRMs and CREs.
4.1. The hyperfinite dictionary
Working on the standing Loeb model $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ from Notation 3.2, we may regard elements of $L^{\infty}$ as essentially bounded Loeb-measurable functions on $I_{N}$. The Loeb integration theory provides a canonical internal representation of such functions.
Lemma 4.1 (Internal liftings and Loeb expectation).
[…]
$\int_{I_{N}}X\,dL(\mu_{N})=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}\tilde{X}(k)\Big).$
If $Y$ is Loeb integrable and $\tilde{Y}$ is an $S$-integrable internal lifting, then
$\int_{I_{N}}Y\,dL(\mu_{N})=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}\tilde{Y}(k)\Big)$
as well.
Proof (reference).
See [10, Chs. 4–5] for liftings of Loeb-measurable functions and the identification of Loeb integrals with standard parts of internal counting integrals. ∎
The following lemma ensures that uniform integrability passes to the non-standard extension, which is crucial for making our hyperfinite representation internal.
Lemma 4.2 (Uniform integrability implies $S$-integrability of $^{*}\mathcal{Z}$).
Let $\mathcal{Z}\subset L^{1}_{+}$ be uniformly integrable with $\sup_{Z\in\mathcal{Z}}\mathsf{E}[Z]\leqslant C<\infty$. Then for every $Z\in{}^{*}\mathcal{Z}$, the internal function $Z:I_{N}\to{}^{*}[0,\infty)$ is $S$-integrable on the Loeb space $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$, with
$\int Z\,dL(\mu_{N})=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}Z(k)\Big).$
In particular, if $\mathsf{E}[Z]=1$ for all $Z\in\mathcal{Z}$, then $\frac{1}{N}\sum_{k=1}^{N}Z(k)\approx 1$ for all $Z\in{}^{*}\mathcal{Z}$.
Moreover, if $X\in L^{\infty}$ has a bounded internal lifting $\tilde{X}$ with $|\tilde{X}|\leqslant M$, then for any $Z\in{}^{*}\mathcal{Z}$ the product $\tilde{X}\cdot Z$ is $S$-integrable and
$\mathsf{E}[-XZ]=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big).$
Proof.
Uniform integrability of $\mathcal{Z}$ means: for every $\varepsilon>0$ there exists $K>0$ such that $\sup_{Z\in\mathcal{Z}}\mathsf{E}[Z\mathbf{1}_{\{Z>K\}}]<\varepsilon$. This is a first-order statement:
$\forall\varepsilon\in\mathbb{Q}_{>0}\ \exists K\in\mathbb{Q}_{>0}\ \forall Z\in\mathcal{Z}:\quad\mathsf{E}[Z\mathbf{1}_{\{Z>K\}}]<\varepsilon.$
By transfer, for every standard $\varepsilon>0$ there exists standard $K>0$ such that $^{*}\mathsf{E}[Z\mathbf{1}_{\{Z>K\}}]<\varepsilon$ for all $Z\in{}^{*}\mathcal{Z}$.
In terms of the hyperfinite counting measure, this says
$\frac{1}{N}\sum_{k:Z(k)>K}Z(k)<\varepsilon$
for all $Z\in{}^{*}\mathcal{Z}$. This is exactly the criterion for $S$-integrability: the “tail” of the internal sum is uniformly small for large standard $K$. By the standard characterisation of $S$-integrability (see [3, Prop. 4.3.5]), each $Z\in{}^{*}\mathcal{Z}$ is $S$-integrable.
The Loeb integral identity follows from the definition of $S$-integrability: if $Z$ is $S$-integrable, then $\int Z\,dL(\mu_{N})=\operatorname{st}(\frac{1}{N}\sum_{k=1}^{N}Z(k))$.
For the constraint $\mathsf{E}[Z]=1$, transfer gives $\int Z\,dL(\mu_{N})=1$ for all $Z\in{}^{*}\mathcal{Z}$. By the Loeb integral identity, $\operatorname{st}(\frac{1}{N}\sum_{k=1}^{N}Z(k))=1$, hence $\frac{1}{N}\sum_{k=1}^{N}Z(k)\approx 1$ for all $Z\in{}^{*}\mathcal{Z}$.
For the product: since $|\tilde{X}|\leqslant M$ and $Z\geqslant 0$ is $S$-integrable, $|\tilde{X}\cdot Z|\leqslant M\cdot Z$ is dominated by an $S$-integrable function, hence $\tilde{X}\cdot Z$ is $S$-integrable. ∎
4.2. Hyperfinite robust representation
We can now state and prove the hyperfinite version of the robust representation theorem. The key observation is that standard attainment on a compact set immediately gives us an internal maximiser via the non-standard extension of the standard maximiser. By Lemma 4.2, elements of $^{*}\mathcal{Z}$ are automatically $S$-integrable, so we can work directly with $^{*}\mathcal{Z}$ as an internal set.
Theorem 4.3 (Hyperfinite robust representation on $L^{\infty}$).
Let $(\Omega,\mathscr{G},\mathsf{P})$ be atomless and separable, and let $\rho:L^{\infty}\to\mathbb{R}$ be a coherent risk measure with the Fatou property. Assume it admits a robust representation of the form
$\rho(X)=\sup_{Z\in\mathcal{Z}}\mathsf{E}[-XZ],\qquad\mathcal{Z}\subset L^{1}_{+},\ \mathsf{E}[Z]=1,$
where $\mathcal{Z}$ is compact in $\sigma(L^{1},L^{\infty})$ (equivalently, uniformly integrable and $\sigma(L^{1},L^{\infty})$-closed).
We may (and do) identify the underlying atomless probability space with the standing Loeb model $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$ (Notation 3.2), and let $\tilde{X}$ be a bounded internal lifting of $X$ as in Lemma 4.1. Then
$\rho(X)=\operatorname{st}\Big(\sup_{Z\in{}^{*}\mathcal{Z}}\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big),\qquad X\in L^{\infty},$
where $^{*}\mathcal{Z}$ is the internal non-standard extension of $\mathcal{Z}$, and elements $Z\in{}^{*}\mathcal{Z}$ are $S$-integrable by Lemma 4.2. The supremum over $^{*}\mathcal{Z}$ is internal, and is attained by some $Z^{\sharp}\in{}^{*}\mathcal{Z}$.
Equivalently, defining weight vectors $a=\Psi(Z)$ via the normalisation
$\Psi(Z)_{k}:=Z(k)/\sum_{j=1}^{N}Z(j),$
we have $\sum_{k}a_{k}=1$ exactly and
$\rho(X)=\operatorname{st}\Big(\sup_{a\in\mathcal{A}_{N}}\sum_{k=1}^{N}a_{k}\,(-\tilde{X}(k))\Big),$
where $\mathcal{A}_{N}:=\{\Psi(Z):Z\in{}^{*}\mathcal{Z}\}$ is internal (as the image of an internal set under an internal map). For the maximiser $Z^{\sharp}\in{}^{*}\mathcal{Z}$, $\Psi(Z^{\sharp})$ achieves a value infinitely close to the supremum over $\mathcal{A}_{N}$.
Proof.
Internality of $^{*}\mathcal{Z}$. Fix once and for all a measurable representative for each element of $\mathcal{Z}\subset L^{1}_{+}$ (possible since $\mathcal{Z}$ is a set, not a quotient). With this choice, $^{*}\mathcal{Z}$ may be viewed as an internal family of internal functions $I_{N}\to{}^{*}[0,\infty)$.
By Lemma 4.2, since $\mathcal{Z}$ is uniformly integrable, every $Z\in{}^{*}\mathcal{Z}$ is automatically $S$-integrable. Crucially, $^{*}\mathcal{Z}$ is an internal set (as the non-standard extension of a standard set), so the supremum $\sup_{Z\in{}^{*}\mathcal{Z}}(\cdots)$ is internal.
Fix $X\in L^{\infty}$ and define the functional $f_{X}:\mathcal{Z}\to\mathbb{R}$ by $f_{X}(Z):=\mathsf{E}[-XZ]$. This is continuous in the $\sigma(L^{1},L^{\infty})$ topology on $\mathcal{Z}$ (by definition of that topology). Since $\mathcal{Z}$ is $\sigma(L^{1},L^{\infty})$-compact, there exists a standard maximiser $Z^{*}\in\mathcal{Z}$ such that $f_{X}(Z^{*})=\rho(X)$.
The non-standard extension $^{*}Z^{*}\in{}^{*}\mathcal{Z}$ (viewing $Z^{*}$ as an element of the standard universe) satisfies $^{*}f_{X}({}^{*}Z^{*})={}^{*}\rho(X)=\rho(X)$ by transfer. Set $Z^{\sharp}:={}^{*}Z^{*}$. Alternatively, one may transfer the existence statement directly: the first-order statement $\exists Z\in\mathcal{Z}\ (f_{X}(Z)=\rho(X))$” holds, hence by transfer there exists $Z^{\sharp}\in{}^{*}\mathcal{Z}$ such that $^{*}f_{X}(Z^{\sharp})=\rho(X)$.
We may (and do) identify the underlying atomless probability space with the standing Loeb model $(I_{N},\mathscr{I}_{N}^{L},L(\mu_{N}))$, and let $\tilde{X}:I_{N}\to{}^{*}\mathbb{R}$ be a bounded internal lifting of $X$ as in Lemma 4.1. By Lemma 4.2, since $Z^{\sharp}\in{}^{*}\mathcal{Z}$ is $S$-integrable and $\tilde{X}$ is bounded, the product $(-\tilde{X})\cdot Z^{\sharp}$ is $S$-integrable. Applying Lemma 4.2 yields
(12) $\rho(X)={}^{*}f_{X}(Z^{\sharp})=\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z^{\sharp}(k)\Big).$
Moreover, by transfer of $f_{X}(Z)\leqslant\rho(X)$ for all $Z\in\mathcal{Z}$”, we have $^{*}f_{X}(Z)\leqslant\rho(X)$ for all $Z\in{}^{*}\mathcal{Z}$. Since the supremum over the internal set $^{*}\mathcal{Z}$ is internal, we obtain
$\rho(X)=\operatorname{st}\Big(\sup_{Z\in{}^{*}\mathcal{Z}}\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big).$
For the normalised weight formulation, define
$H(Z):=\frac{1}{N}\sum_{j=1}^{N}Z(j),\qquad\Psi(Z)_{k}:=\frac{Z(k)}{\sum_{j=1}^{N}Z(j)},\quad k=1,\dots,N.$
Then $\sum_{k=1}^{N}\Psi(Z)_{k}=1$ exactly (this is an algebraic identity).
For $Z\in{}^{*}\mathcal{Z}$, by Lemma 4.2 we have $\frac{1}{N}\sum_{k=1}^{N}Z(k)\approx 1$, i.e., $H(Z)\approx 1$. Since $Z\geqslant 0$ and $H(Z)\approx 1>0$, $\Psi(Z)$ is well-defined and internal. The set $\mathcal{A}_{N}:=\{\Psi(Z):Z\in{}^{*}\mathcal{Z}\}$ is internal (as the image of an internal set under an internal map).
The connection between the two formulations is:
$\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)=H(Z)\sum_{k=1}^{N}\Psi(Z)_{k}\,(-\tilde{X}(k)).$
Since $|\tilde{X}|\leqslant M$ for some standard $M$, the sum $\sum_{k=1}^{N}\Psi(Z)_{k}(-\tilde{X}(k))$ is bounded (in absolute value by $M$, since $\sum_{k}\Psi(Z)_{k}=1$). Since $H(Z)\approx 1$,
$\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big)=\operatorname{st}\Big(\sum_{k=1}^{N}\Psi(Z)_{k}\,(-\tilde{X}(k))\Big).$
The weight-vector supremum satisfies
$\operatorname{st}\Big(\sup_{a\in\mathcal{A}_{N}}\sum_{k=1}^{N}a_{k}\,(-\tilde{X}(k))\Big)=\operatorname{st}\Big(\sup_{Z\in{}^{*}\mathcal{Z}}\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))\,Z(k)\Big).$
Indeed, for $Z\in{}^{*}\mathcal{Z}$ we have
$\sum_{k=1}^{N}\Psi(Z)(k)(-\tilde{X}(k))=\frac{1}{\sum_{j=1}^{N}Z(j)}\sum_{k=1}^{N}(-\tilde{X}(k))Z(k)=\frac{1}{H(Z)}\cdot\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))Z(k).$
Since $^{*}\mathcal{Z}$ is internal and $H(Z)\approx 1$ for every $Z\in{}^{*}\mathcal{Z}$, the internal set $\{|1/H(Z)-1|:Z\in{}^{*}\mathcal{Z}\}$ consists only of infinitesimals, hence its supremum is infinitesimal. Moreover,
$\left|\frac{1}{N}\sum_{k=1}^{N}(-\tilde{X}(k))Z(k)\right|\leqslant\|\tilde{X}\|_{\infty}\,H(Z),$
so taking internal suprema preserves the standard part. Therefore the two internal suprema have the same standard part.
The attainment statement follows from transfer: the standard maximiser $Z^{*}\in\mathcal{Z}$ extends to $Z^{\sharp}={}^{*}Z^{*}\in{}^{*}\mathcal{Z}$, and $a^{\sharp}:=\Psi(Z^{\sharp})$ achieves a value infinitely close to the supremum over $\mathcal{A}_{N}$. ∎
Remark 4.4 (Why compactness appears).
The passage from a standard supremum $\sup_{Q\in\mathcal{Q}}\mathsf{E}_{Q}[-X]$ to an internal maximum over $^{*}\mathcal{Q}$ uses the NSA characterisation of compactness. In the absence of compactness one still obtains $\varepsilon$-maximisers in $^{*}\mathcal{Q}$ and hence a hyperfinite representation up to infinitesimals, uniformly over $X$ with $\|X\|_{\infty}\leqslant 1$; compactness is the natural condition ensuring exact shadowing by standard parts.
Remark 4.5 (The hyperfinite dictionary).
Theorem 4.3 establishes the central dictionary of our approach:
CRM (population) Hyperfinite representation
Probability measure $Q$ Weight vector $a\in({}^{*}[0,1])^{N}$, $\sum a_{k}=1$
Expectation $\mathsf{E}_{Q}[-X]$ Hyperfinite sum $\sum_{k=1}^{N}a_{k}(-x_{k})$
Supremum over $\mathcal{Q}$ Internal supremum over $\mathcal{A}_{N}$
Risk measure $\rho(X)$ Standard part of hyperfinite support function
4.3. From hyperfinite representation to CRE representation
The finite-sample representation theorems for CREs now emerge as special cases.
Hyperfinite proof of Theorem 2.9.
[…]
A CRE $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ induces a CRM $\rho$ on this finite space by $\rho(X):=\hat{\rho}_{n}(x)$ where $x_{i}=X(\omega_{i})$. The axioms (E1)–(E4) translate directly to (R1)–(R4).
By finite-dimensional convex duality (the finite space version of the robust representation, which does not require atomlessness), there exists a convex set $\mathcal{Q}$ of probability measures on $\hat{\Omega}$ such that
$\rho(X)=\sup_{Q\in\mathcal{Q}}\mathsf{E}_{Q}[-X].$
Each $Q$ on $\hat{\Omega}$ corresponds to a weight vector $a\in\Delta_{n}$ via $a_{i}=Q(\{\omega_{i}\})$, in which case $\mathsf{E}_{Q}[-X]=\sum_{i=1}^{n}a_{i}(-x_{i})=\langle a,-x\rangle$. Setting $M^{*}_{\hat{\rho}_{n}}:=\{a(Q):Q\in\mathcal{Q}\}$, we obtain
$\hat{\rho}_{n}(x)=\sup_{a\in M^{*}_{\hat{\rho}_{n}}}\langle a,-x\rangle.$
Convexity of $M^{*}_{\hat{\rho}_{n}}$ follows from convexity of $\mathcal{Q}$, and attainment follows from compactness of the simplex $\Delta_{n}$. ∎
For the law-invariant representation, we need the rearrangement inequality.
Lemma 4.6 (Rearrangement inequality [12, Theorem 368]).
Let $u,v\in\mathbb{R}^{n}$ with decreasing rearrangements $u^{\downarrow}$ and $v^{\downarrow}$. Then for any permutation $\sigma$,
$\sum_{i=1}^{n}u_{i}v_{\sigma(i)}\leqslant\sum_{i=1}^{n}u^{\downarrow}_{i}v^{\downarrow}_{i},$
with equality when $u$ and $v$ are both sorted in the same order (both increasing or both decreasing).
We may now prove Theorem 2.10.
Proof of Theorem 2.10.
By Theorem 2.9 there exists a non-empty convex $M^{*}_{\hat{\rho}_{n}}\subseteq\Delta_{n}$ such that $\hat{\rho}_{n}(x)=\sup_{a\in M^{*}_{\hat{\rho}_{n}}}\langle a,-x\rangle$.
Define the symmetrised set
$\widetilde{M}:=\operatorname{conv}\bigl\{\pi(a):\ a\in M^{*}_{\hat{\rho}_{n}},\ \pi\ \text{a permutation of }\{1,\dots,n\}\bigr\}\subseteq\Delta_{n},$
where $\pi(a)$ denotes the permuted vector $(a_{\pi(1)},\dots,a_{\pi(n)})$. Then $\widetilde{M}$ is convex and non-empty. Moreover, law invariance implies that for every $x$,
$\hat{\rho}_{n}(x)=\hat{\rho}_{n}(\pi(x))=\sup_{a\in M^{*}_{\hat{\rho}_{n}}}\langle a,-\pi(x)\rangle=\sup_{a\in M^{*}_{\hat{\rho}_{n}}}\langle\pi^{-1}(a),-x\rangle,$
and taking the supremum over all permutations shows
$\hat{\rho}_{n}(x)=\sup_{b\in\widetilde{M}}\langle b,-x\rangle.$
Now for each $b\in\Delta_{n}$ let $b^{\downarrow}\in\Delta_{n}^{\downarrow}$ be its decreasing rearrangement. Note that s(x)-s(x) is the decreasing rearrangement of x-x (since $s(x)=(x_{1:n},\ldots,x_{n:n})$ is increasing). By the rearrangement inequality (Lemma 4.6) applied to $u=b$ and v=xv=-x,
$\langle b,-x\rangle=\sum_{i=1}^{n}b_{i}(-x_{i})\leqslant\sum_{i=1}^{n}b^{\downarrow}_{i}(-x)^{\downarrow}_{i}=\sum_{i=1}^{n}b^{\downarrow}_{i}(-x_{i:n})=\langle b^{\downarrow},-s(x)\rangle.$
Hence
$\hat{\rho}_{n}(x)=\sup_{b\in\widetilde{M}}\langle b,-x\rangle\leqslant\sup_{b\in\widetilde{M}}\langle b^{\downarrow},-s(x)\rangle.$
Conversely, for any $b\in\widetilde{M}$, since $\widetilde{M}$ is permutation invariant, we have $b^{\downarrow}\in\widetilde{M}$. Moreover, $s(x)$ is a permutation of $x$, so there exists a permutation $\sigma$ with $s(x)=\sigma(x)$. Then
$\langle b^{\downarrow},-s(x)\rangle=\langle\sigma^{-1}(b^{\downarrow}),-x\rangle\leqslant\sup_{b^{\prime}\in\widetilde{M}}\langle b^{\prime},-x\rangle=\hat{\rho}_{n}(x),$
where we used $\sigma^{-1}(b^{\downarrow})\in\widetilde{M}$ by permutation invariance. Hence $\sup_{b\in\widetilde{M}}\langle b^{\downarrow},-s(x)\rangle\leqslant\hat{\rho}_{n}(x)$, establishing equality.
Therefore, with
$M^{s}_{\hat{\rho}_{n}}:=\operatorname{cl}\operatorname{conv}\{b^{\downarrow}:\ b\in\widetilde{M}\}\subseteq\Delta_{n}^{\downarrow},$
we obtain
$\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i:n}).$
Attainment follows because $M^{s}_{\hat{\rho}_{n}}$ is a closed subset of the compact simplex $\Delta_{n}^{\downarrow}$, and the objective is continuous and linear in $a$. ∎
Proof of Theorem 2.11 (finite Choquet / distortion representation).
Let $\hat{\rho}_{n}$ be coherent, law-invariant, and comonotonic additive. By Theorem 2.10 there exists a non-empty convex set $M^{s}_{\hat{\rho}_{n}}\subseteq\Delta_{n}^{\downarrow}$ such that
$\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i:n}).$
A standard representation theorem for comonotonic additive, monotone functionals on a finite lattice states that such a functional is a Choquet integral with respect to a unique capacity (called the core of the functional); see, e.g., [9, Ch. 5]. In the present finite-dimensional setting, one may take the capacity $c:2^{\{1,\dots,n\}}\to[0,1]$ defined by
$c(A):=\hat{\rho}_{n}(-\mathbf{1}_{A}),\qquad A\subseteq\{1,\dots,n\},$
where $\mathbf{1}_{A}\in\mathbb{R}^{n}$ is the indicator vector of $A$.
Law invariance implies that $c(A)$ depends only on $|A|$, hence $c$ is determined by a unique distortion function on the grid $\{0,1,\dots,n\}$:
$g(k):=c(\{1,2,\dots,k\})=\hat{\rho}_{n}(-\mathbf{1}_{\{1,\dots,k\}}),\qquad k=0,\dots,n,$
with $g(0)=0$ and $g(n)=1$. Coherence (in particular subadditivity) is equivalent in this finite setting to discrete concavity of $g$ (equivalently, to $\Delta_{n}^{\downarrow}$-monotonicity of the resulting weights).
By the finite Choquet integral formula (see, e.g., [9, Ch. 5] or the risk-measure presentation in [11]), for every $x\in\mathbb{R}^{n}$,
$\hat{\rho}_{n}(x)=\sum_{i=1}^{n}\bigl(g(i)-g(i-1)\bigr)\,(-x_{i:n}).$
Setting $a_{i}:=g(i)-g(i-1)$ yields $a\in\Delta_{n}^{\downarrow}$. Uniqueness follows because the values $\hat{\rho}_{n}(-\mathbf{1}_{\{1,\dots,k\}})=g(k)$ recover the increments $a_{i}=g(i)-g(i-1)$, hence the weight vector is uniquely determined.
Consequently $M^{s}_{\hat{\rho}_{n}}$ must be the singleton $\{a\}$, proving the theorem. ∎
5. Discrete Kusuoka representation for law-invariant CREs
The Kusuoka representation (Theorem 2.5) expresses law-invariant CRMs as suprema over mixtures of expected shortfall. In this section, we establish the finite-sample analogue for CREs.
5.1. Discrete expected shortfall
We begin by defining the natural finite-sample version of expected shortfall.
Definition 5.1 (Discrete expected shortfall).
For $x\in\mathbb{R}^{n}$ and $k\in\{1,\ldots,n\}$, the discrete expected shortfall at level $k/n$ is
(13) $\operatorname{dES}_{k/n}(x):=-\frac{1}{k}\sum_{i=1}^{k}x_{i:n}.$
This is the average of the $k$ smallest (most negative) outcomes, negated to give a positive quantity for losses. When $k=1$, we recover the minimum: $\operatorname{dES}_{1/n}(x)=-x_{1:n}$. When $k=n$, we get the negative mean: $\operatorname{dES}_{1}(x)=-\frac{1}{n}\sum_{i=1}^{n}x_{i}$.
Remark 5.2.
The discrete expected shortfall $\operatorname{dES}_{k/n}$ is a coherent, law-invariant, comonotonic risk estimator. Its weight vector is $a=(\frac{1}{k},\ldots,\frac{1}{k},0,\ldots,0)$ with $k$ entries equal to $1/k$, which lies in $\Delta_{n}^{\downarrow}$.
5.2. Decomposition of L-estimators into discrete expected shortfall
The key technical result is that any L-estimator with non-increasing weights can be written as a mixture of discrete expected shortfalls.
Lemma 5.3 (ES-decomposition lemma).
Let $a=(a_{1},\ldots,a_{n})\in\Delta_{n}^{\downarrow}$ and set $a_{n+1}:=0$. Define
(14) $\mu_{k}:=k(a_{k}-a_{k+1}),\qquad k=1,\ldots,n.$
Then:
  1. (i)
    $\mu_{k}\geqslant 0$ for all $k$;
  2. (ii)
    $\sum_{k=1}^{n}\mu_{k}=1$;
  3. (iii)
    for every $x\in\mathbb{R}^{n}$,
    (15) $\sum_{i=1}^{n}a_{i}(-x_{i:n})=\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x).$
Proof.
(i) Since $a$ is non-increasing, $a_{k}\geqslant a_{k+1}$, hence $\mu_{k}=k(a_{k}-a_{k+1})\geqslant 0$.
(ii) We compute:
$\sum_{k=1}^{n}\mu_{k}$ $=\sum_{k=1}^{n}k(a_{k}-a_{k+1})$
$=\sum_{k=1}^{n}ka_{k}-\sum_{k=1}^{n}ka_{k+1}$
$=\sum_{k=1}^{n}ka_{k}-\sum_{j=2}^{n+1}(j-1)a_{j}$
$=a_{1}+\sum_{k=2}^{n}ka_{k}-\sum_{j=2}^{n}(j-1)a_{j}-na_{n+1}$
$=a_{1}+\sum_{k=2}^{n}(k-(k-1))a_{k}-0$
$=a_{1}+\sum_{k=2}^{n}a_{k}=\sum_{k=1}^{n}a_{k}=1.$
(iii) We expand the right-hand side:
$\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x)$ $=\sum_{k=1}^{n}\mu_{k}\cdot\left(-\frac{1}{k}\sum_{i=1}^{k}x_{i:n}\right)$
$=-\sum_{k=1}^{n}(a_{k}-a_{k+1})\sum_{i=1}^{k}x_{i:n}.$
We exchange the order of summation. For each $i\in\{1,\ldots,n\}$, the term $x_{i:n}$ appears in the inner sum for all $k\geqslant i$. Thus:
$-\sum_{k=1}^{n}(a_{k}-a_{k+1})\sum_{i=1}^{k}x_{i:n}$ $=-\sum_{i=1}^{n}x_{i:n}\sum_{k=i}^{n}(a_{k}-a_{k+1})$
$=-\sum_{i=1}^{n}x_{i:n}(a_{i}-a_{n+1})$
$=-\sum_{i=1}^{n}x_{i:n}\cdot a_{i}$
$=\sum_{i=1}^{n}a_{i}(-x_{i:n}).\qed$
5.3. The discrete Kusuoka representation theorem
Let $\mathcal{T}:\Delta_{n}^{\downarrow}\to\Delta_{n}$ denote the linear map $a\mapsto\mu$ defined by (14). The lemma shows that $\mathcal{T}$ is well-defined and maps non-increasing probability vectors to probability vectors.
Remark 5.4 (Inverse of $\mathcal{T}$).
Given $\mu\in\Delta_{n}$ with $\mu_{k}=k(a_{k}-a_{k+1})$ and $a_{n+1}=0$, one recovers
$a_{i}=\sum_{k=i}^{n}\frac{\mu_{k}}{k},\qquad i=1,\ldots,n.$
Thus mixtures of $\operatorname{dES}_{k/n}$ are in one-to-one correspondence with weight vectors in $\Delta_{n}^{\downarrow}$.
Theorem 5.5 (Discrete Kusuoka representation).
Let $\hat{\rho}_{n}:\mathbb{R}^{n}\to\mathbb{R}$ be a law-invariant CRE. Then there exists a non-empty convex set $\mathcal{M}_{n}\subseteq\Delta_{n}$ such that for all $x\in\mathbb{R}^{n}$,
(16) $\hat{\rho}_{n}(x)=\sup_{\mu\in\mathcal{M}_{n}}\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x),$
with the supremum attained. Moreover, one may take $\mathcal{M}_{n}=\mathcal{T}(M^{s}_{\hat{\rho}_{n}})$, where $M^{s}_{\hat{\rho}_{n}}$ is the representing set from Theorem 2.10.
Proof.
By Theorem 2.10, there exists a convex set $M^{s}_{\hat{\rho}_{n}}\subseteq\Delta_{n}^{\downarrow}$ such that
$\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{i=1}^{n}a_{i}(-x_{i:n}).$
By Lemma 5.3, for each $a\in M^{s}_{\hat{\rho}_{n}}$,
$\sum_{i=1}^{n}a_{i}(-x_{i:n})=\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x),$
where $\mu=\mathcal{T}(a)$. Thus
$\hat{\rho}_{n}(x)=\sup_{a\in M^{s}_{\hat{\rho}_{n}}}\sum_{k=1}^{n}\mathcal{T}(a)_{k}\,\operatorname{dES}_{k/n}(x)=\sup_{\mu\in\mathcal{T}(M^{s}_{\hat{\rho}_{n}})}\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x).$
Setting $\mathcal{M}_{n}:=\mathcal{T}(M^{s}_{\hat{\rho}_{n}})$, we obtain (16).
Convexity: since $\mathcal{T}$ is linear, $\mathcal{T}(M^{s}_{\hat{\rho}_{n}})$ is convex whenever $M^{s}_{\hat{\rho}_{n}}$ is.
Attainment: the supremum in (6) is attained for each $x$ at some $a^{*}\in M^{s}_{\hat{\rho}_{n}}$. Then $\mu^{*}=\mathcal{T}(a^{*})$ achieves the supremum in (16). ∎
Corollary 5.6 (Comonotonic case).
A CRE $\hat{\rho}_{n}$ is comonotonic and law-invariant if and only if there exists a unique $\mu\in\Delta_{n}$ such that
$\hat{\rho}_{n}(x)=\sum_{k=1}^{n}\mu_{k}\,\operatorname{dES}_{k/n}(x),\qquad x\in\mathbb{R}^{n}.$
Proof.
By Theorem 2.11, $M^{s}_{\hat{\rho}_{n}}=\{a\}$ is a singleton, hence $\mathcal{M}_{n}=\{\mathcal{T}(a)\}$ is a singleton. ∎
Remark 5.7.
The discrete Kusuoka representation (16) is the finite-sample analogue of the population-level Kusuoka representation (3). The probability measure $\nu$ on $(0,1]$ is replaced by a probability vector $\mu\in\Delta_{n}$, and the continuous expected shortfall $\operatorname{ES}_{\alpha}$ is replaced by its discrete counterpart $\operatorname{dES}_{k/n}$. This representation makes explicit the sense in which CREs are “statistical shadows” of CRMs.
6. Spectral risk measures and hyperfinite L-statistics
We now turn to spectral risk measures and their finite-sample estimators, developing the hyperfinite perspective that will underpin our consistency results.
6.1. Spectral risk as a hyperfinite L-statistic
Let $\varphi$ be a spectrum (Definition 2.6) and $X\in L^{1}$ a random variable with lower quantile function $q_{X}$. The spectral risk measure is
$\rho_{\varphi}(X)=-\int_{0}^{1}q_{X}(\alpha)\,\varphi(\alpha)\,d\alpha.$
Now consider a hyperfinite i.i.d. sample $(X_{1},\ldots,X_{N})$ with $N\in{}^{*}\mathbb{N}$ infinite. Let $X_{1:N}\leqslant\cdots\leqslant X_{N:N}$ be the order statistics and $\alpha_{k}:=k/N$. Define the hyperfinite L-statistic
(17) $L_{N}^{\varphi}:=-\frac{1}{N}\sum_{k=1}^{N}\varphi(\alpha_{k})\,X_{k:N}.$
Proposition 6.1 (Hyperfinite spectral representation).
Let $X\in L^{1}$ with quantile function $q_{X}$, and let $\varphi$ be a bounded spectrum. Then, for Loeb-almost all sample paths,
(18) $\rho_{\varphi}(X)=\operatorname{st}(L_{N}^{\varphi}).$
Proof.
Apply Theorem 3.20 with $g=\varphi$. Since $\varphi$ is bounded and (being monotone) Riemann integrable, and $X\in L^{1}$ implies $\int_{0}^{1}|\varphi(\alpha)q_{X}(\alpha)|\,d\alpha<\infty$, we obtain
$\operatorname{st}\Big(\frac{1}{N}\sum_{k=1}^{N}\varphi(\alpha_{k})X_{k:N}\Big)=\int_{0}^{1}\varphi(\alpha)q_{X}(\alpha)\,d\alpha.$
Multiplying by 1-1 gives $\operatorname{st}(L_{N}^{\varphi})=\rho_{\varphi}(X)$. ∎
6.2. Canonical spectral plug-in estimators
Given a spectrum $\varphi$ and sample size $n\in\mathbb{N}$, the canonical finite-sample estimator is constructed as follows.
Definition 6.2 (Canonical spectral plug-in estimator).
For $x\in\mathbb{R}^{n}$, define
(19) $\hat{\rho}_{n,\varphi}(x):=-\sum_{i=1}^{n}a_{i,n}(\varphi)\,x_{i:n},$
where the weights are
(20) $a_{i,n}(\varphi):=\int_{(i-1)/n}^{i/n}\varphi(s)\,ds,\qquad i=1,\ldots,n.$
Proposition 6.3.
The canonical spectral plug-in estimator $\hat{\rho}_{n,\varphi}$ is a comonotonic, law-invariant CRE. Moreover, its Kusuoka representation (Theorem 5.5) has $\mathcal{M}_{n}=\{\mu\}$ a singleton with $\mu=\mathcal{T}(a_{n}(\varphi))$.
Proof.
Since $\varphi$ is non-increasing, the weights $a_{i,n}(\varphi)$ are non-increasing in $i$ (as integrals of a non-increasing function over consecutive intervals). Also, $\sum_{i=1}^{n}a_{i,n}(\varphi)=\int_{0}^{1}\varphi=1$, so $a_{n}(\varphi)\in\Delta_{n}^{\downarrow}$.
By Theorem 2.11, $\hat{\rho}_{n,\varphi}$ is a comonotonic, law-invariant CRE. The Kusuoka representation follows from Corollary 5.6. ∎
6.3. Step approximations and the spectral-estimator correspondence
For theoretical analysis, it is useful to associate with each weight vector $a_{n}\in\Delta_{n}^{\downarrow}$ a step function $\varphi_{n}$ on $[0,1]$.
Definition 6.4 (Associated step function).
For $a_{n}=(a_{1,n},\ldots,a_{n,n})\in\Delta_{n}^{\downarrow}$, define
(21) $\varphi_{n}(t):=\sum_{i=1}^{n}na_{i,n}\,\mathbf{1}_{((i-1)/n,i/n]}(t),\qquad t\in(0,1].$
When $a_{n}$ arises from a spectrum $\varphi$ via (20), we have
$\varphi_{n}(t)=n\int_{(i-1)/n}^{i/n}\varphi(s)ds\text{ for }t\in((i-1)/n,i/n],$
which is the average of $\varphi$ on that interval.
Lemma 6.5 (Integral correspondence).
For $a_{n}\in\Delta_{n}^{\downarrow}$ and its associated step function $\varphi_{n}$,
(22) $\sum_{i=1}^{n}a_{i,n}\,x_{i:n}=\int_{0}^{1}\varphi_{n}(\alpha)\,q_{n}(\alpha)\,d\alpha,$
where $q_{n}(\alpha):=x_{\lceil n\alpha\rceil:n}$ is the empirical quantile function.
Proof.
For $\alpha\in((i-1)/n,i/n]$, we have $\lceil n\alpha\rceil=i$, so $q_{n}(\alpha)=x_{i:n}$. Thus
$\int_{0}^{1}\varphi_{n}(\alpha)q_{n}(\alpha)d\alpha=\sum_{i=1}^{n}\int_{(i-1)/n}^{i/n}na_{i,n}\cdot x_{i:n}\,d\alpha=\sum_{i=1}^{n}na_{i,n}\cdot\frac{1}{n}\cdot x_{i:n}=\sum_{i=1}^{n}a_{i,n}x_{i:n}.\qed$
7. Spectral plug-in consistency
We now establish the consistency of spectral plug-in estimators, culminating in a uniform consistency theorem over Lipschitz families of spectra.
7.1. Pointwise consistency: analytic core
The following lemma isolates the key analytic condition for consistency.
Lemma 7.1 (Primitive convergence implies density convergence).
Let $(\varphi_{n})_{n\in\mathbb{N}}$ be a sequence of step functions on $[0,1]$, each non-increasing, non-negative, and satisfying $\int_{0}^{1}\varphi_{n}=1$. Assume $\sup_{n}\|\varphi_{n}\|_{\infty}<\infty$. Let $\varphi$ be a bounded spectrum. Define the primitives
$\Phi_{n}(t):=\int_{0}^{t}\varphi_{n}(s)\,ds,\qquad\Phi(t):=\int_{0}^{t}\varphi(s)\,ds.$
If $\Phi_{n}(t)\to\Phi(t)$ for all $t\in(0,1)$, then:
  1. (i)
    $\varphi_{n}(t)\to\varphi(t)$ for almost every $t\in(0,1)$;
  2. (ii)
    $\varphi_{n}\to\varphi$ in $L^{1}([0,1])$.
Proof.
(i) Fix $t\in(0,1)$ and $h>0$ sufficiently small. Since $\varphi_{n}$ is non-increasing, for any $s\in[t,t+h]$ we have $\varphi_{n}(s)\leqslant\varphi_{n}(t)$, hence
$\frac{\Phi_{n}(t+h)-\Phi_{n}(t)}{h}=\frac{1}{h}\int_{t}^{t+h}\varphi_{n}(s)ds\leqslant\varphi_{n}(t).$
Similarly, for $s\in[t-h,t]$, $\varphi_{n}(s)\geqslant\varphi_{n}(t)$, so
$\varphi_{n}(t)\leqslant\frac{\Phi_{n}(t)-\Phi_{n}(t-h)}{h}.$
Taking $\liminf$ and $\limsup$ as $n\to\infty$, and using $\Phi_{n}\to\Phi$ pointwise:
$\frac{\Phi(t+h)-\Phi(t)}{h}\leqslant\liminf_{n\to\infty}\varphi_{n}(t)\leqslant\limsup_{n\to\infty}\varphi_{n}(t)\leqslant\frac{\Phi(t)-\Phi(t-h)}{h}.$
Now let $h\downarrow 0$. Since $\varphi$ is non-increasing, $\Phi$ is concave, hence differentiable almost everywhere. At differentiability points of $\Phi$, both bounds converge to $\Phi^{\prime}(t)=\varphi(t)$, giving $\varphi_{n}(t)\to\varphi(t)$.
(ii) By (i), $\varphi_{n}\to\varphi$ pointwise almost everywhere. The uniform bound $\sup_{n}\|\varphi_{n}\|_{\infty}<\infty$ allows dominated convergence: $\int_{0}^{1}|\varphi_{n}-\varphi|\to 0$. ∎
7.2. Spectral L-estimator consistency theorem
Theorem 7.2 (Spectral L-estimator consistency: a distribution-free criterion).
Let $\varphi$ be a bounded spectrum. Let $\hat{\rho}_{n}(x)=-\sum_{i=1}^{n}a_{i,n}x_{i:n}$ with $a_{n}\in\Delta_{n}^{\downarrow}$, and let $\varphi_{n}$ be the associated step function (21). Assume $\sup_{n}\|\varphi_{n}\|_{\infty}<\infty$.
Then the following are equivalent:
  1. (i)
    For every $t\in(0,1)$, $\int_{0}^{t}\varphi_{n}(s)\,ds\to\int_{0}^{t}\varphi(s)\,ds$.
  2. (ii)
    For every $X\in L^{1}$ and every i.i.d. sample $(X_{i})$ with law $X$,
    $\hat{\rho}_{n}(X_{1},\ldots,X_{n})\xrightarrow[n\to\infty]{\mathrm{a.s.}}\rho_{\varphi}(X).$
Remark 7.3.
The condition $\sup_{n}\|\varphi_{n}\|_{\infty}<\infty$ is automatic for the canonical discretisation $a_{i,n}=\int_{(i-1)/n}^{i/n}\varphi$ of a bounded spectrum $\varphi$: in this case $\|\varphi_{n}\|_{\infty}\leqslant\|\varphi\|_{\infty}$ since $\varphi_{n}$ is a step-function approximation. For general L-estimator weights $(a_{i,n})$ not arising from a fixed spectrum, the uniform bound is a genuine assumption that must be verified.
Proof.
Assume (i). By Lemma 7.1, the primitive convergence implies $\varphi_{n}\to\varphi$ in $L^{1}([0,1])$. Let $q_{n}(\alpha):=X_{\lceil n\alpha\rceil:n}$ be the empirical quantile function and let $q_{X}$ be the population lower quantile. By the Glivenko–Cantelli theorem, $q_{n}(\alpha)\to q_{X}(\alpha)$ almost surely at every continuity point of $q_{X}$, hence for almost every $\alpha\in(0,1)$.
Using Lemma 6.5,
$\hat{\rho}_{n}(X_{1},\ldots,X_{n})=-\int_{0}^{1}\varphi_{n}(\alpha)\,q_{n}(\alpha)\,d\alpha,\qquad\rho_{\varphi}(X)=-\int_{0}^{1}\varphi(\alpha)\,q_{X}(\alpha)\,d\alpha.$
Therefore,
$\left|\hat{\rho}_{n}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right|\leqslant\int_{0}^{1}|\varphi_{n}-\varphi|\,|q_{n}|\,d\alpha+\int_{0}^{1}|\varphi|\,|q_{n}-q_{X}|\,d\alpha.$
We handle these two integrals separately.
Fix $M>0$. Since $|q_{n}|=|q_{n}|\mathbf{1}_{\{|q_{n}|\leqslant M\}}+|q_{n}|\mathbf{1}_{\{|q_{n}|>M\}}$, we obtain
$\int_{0}^{1}|\varphi_{n}-\varphi|\,|q_{n}|\,d\alpha$ $\leqslant M\int_{0}^{1}|\varphi_{n}-\varphi|\,d\alpha+\bigl(\sup_{m}\|\varphi_{m}\|_{\infty}+\|\varphi\|_{\infty}\bigr)\int_{0}^{1}|q_{n}|\,\mathbf{1}_{\{|q_{n}|>M\}}\,d\alpha.$
The first term converges to 0 because $\varphi_{n}\to\varphi$ in $L^{1}$. Moreover,
$\int_{0}^{1}|q_{n}(\alpha)|\,\mathbf{1}_{\{|q_{n}(\alpha)|>M\}}\,d\alpha=\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\,\mathbf{1}_{\{|X_{i}|>M\}},$
since $q_{n}$ is a step function taking the values $X_{i:n}$. By the strong law of large numbers, the right-hand side converges almost surely to $\mathsf{E}[|X|\,\mathbf{1}_{\{|X|>M\}}]$. Letting $n\to\infty$ and then $M\to\infty$ (using $\mathsf{E}|X|<\infty$) shows
$\int_{0}^{1}|\varphi_{n}-\varphi|\,|q_{n}|\,d\alpha\longrightarrow 0\qquad\text{almost surely.}$
For the second term, note first that $q_{n}(\alpha)\to q_{X}(\alpha)$ almost everywhere. Define truncations $q_{n}^{(M)}:=\max(\min(q_{n},M),-M)$ and $q_{X}^{(M)}:=\max(\min(q_{X},M),-M)$. Then $|q_{n}^{(M)}-q_{X}^{(M)}|\leqslant 2M$ and $q_{n}^{(M)}\to q_{X}^{(M)}$ almost everywhere, so dominated convergence yields $\int_{0}^{1}|q_{n}^{(M)}-q_{X}^{(M)}|\,d\alpha\to 0$ almost surely for each fixed $M$. Moreover,
$\int_{0}^{1}|q_{n}-q_{X}|\leqslant\int_{0}^{1}|q_{n}-q_{n}^{(M)}|+\int_{0}^{1}|q_{n}^{(M)}-q_{X}^{(M)}|+\int_{0}^{1}|q_{X}^{(M)}-q_{X}|.$
The middle term tends to 0 almost surely for fixed $M$. The last term tends to 0 as $M\to\infty$ because $q_{X}\in L^{1}(0,1)$ when $X\in L^{1}$. Finally,
$\int_{0}^{1}|q_{n}(\alpha)|\,\mathbf{1}_{\{|q_{n}(\alpha)|>M\}}\,d\alpha=\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\,\mathbf{1}_{\{|X_{i}|>M\}}\longrightarrow\mathsf{E}[|X|\,\mathbf{1}_{\{|X|>M\}}]$
almost surely, and the right-hand side vanishes as $M\to\infty$. Hence $\int_{0}^{1}|q_{n}-q_{X}|\,d\alpha\to 0$ almost surely, and therefore
$\int_{0}^{1}|\varphi(\alpha)|\,|q_{n}(\alpha)-q_{X}(\alpha)|\,d\alpha\leqslant\|\varphi\|_{\infty}\int_{0}^{1}|q_{n}-q_{X}|\longrightarrow 0\qquad\text{almost surely.}$
Combining the two parts proves (ii).
Conversely, assume (ii). Fix $t\in(0,1)$ and let $U\sim\mathrm{Unif}(0,1)$. Set $X:=-\mathbf{1}_{\{U\leqslant t\}}$. Then $X$ is bounded and
$q_{X}(\alpha)=\begin{cases}-1,&\alpha\in(0,t],\\ 0,&\alpha\in(t,1),\end{cases}\qquad\text{so}\qquad\rho_{\varphi}(X)=\int_{0}^{t}\varphi(\alpha)\,d\alpha.$
For an i.i.d. sample $X_{i}:=-\mathbf{1}_{\{U_{i}\leqslant t\}}$, let $m_{n}:=\sum_{i=1}^{n}\mathbf{1}_{\{U_{i}\leqslant t\}}$ and $t_{n}:=m_{n}/n$. The order statistics satisfy $X_{1:n}=\cdots=X_{m_{n}:n}=-1$ and $X_{m_{n}+1:n}=\cdots=X_{n:n}=0$, hence
$\hat{\rho}_{n}(X_{1},\ldots,X_{n})=-\sum_{i=1}^{n}a_{i,n}X_{i:n}=\sum_{i=1}^{m_{n}}a_{i,n}=\int_{0}^{t_{n}}\varphi_{n}(\alpha)\,d\alpha.$
By the strong law, $t_{n}\to t$ almost surely, and the uniform bound $\sup_{n}\|\varphi_{n}\|_{\infty}<\infty$ implies
$\left|\int_{0}^{t_{n}}\varphi_{n}-\int_{0}^{t}\varphi_{n}\right|\leqslant\sup_{n}\|\varphi_{n}\|_{\infty}\,|t_{n}-t|\longrightarrow 0\qquad\text{almost surely.}$
Using (ii) for this bounded $X$ gives $\int_{0}^{t_{n}}\varphi_{n}\to\int_{0}^{t}\varphi$ almost surely, and the previous estimate therefore yields $\int_{0}^{t}\varphi_{n}\to\int_{0}^{t}\varphi$, i.e. (i). ∎
7.3. Uniform consistency over Lipschitz spectral classes
We now establish uniform consistency over families of spectra satisfying Lipschitz and boundedness conditions.
Definition 7.4 (Lipschitz spectral class).
A family $\mathcal{V}$ of spectra is a Lipschitz spectral class with constants $(C,L)$ if:
  1. (V1)
    Each $\varphi\in\mathcal{V}$ is non-increasing, bounded, and $\int_{0}^{1}\varphi=1$.
  2. (V2)
    $\sup_{\varphi\in\mathcal{V}}\|\varphi\|_{\infty}\leqslant C$.
  3. (V3)
    Each $\varphi\in\mathcal{V}$ is $L$-Lipschitz: $|\varphi(s)-\varphi(t)|\leqslant L|s-t|$.
Lemma 7.5 (Uniform discretisation bound).
Let $\mathcal{V}$ be a Lipschitz spectral class with constants $(C,L)$. For each $\varphi\in\mathcal{V}$ and $n\in\mathbb{N}$, let $a_{n}(\varphi)$ be the canonical weights (20) and $\varphi_{n}$ the associated step function. Then:
  1. (i)
    $\sup_{\varphi\in\mathcal{V}}\sup_{t\in(0,1]}|\varphi_{n}(t)-\varphi(t)|\leqslant L/n$.
  2. (ii)
    For any $x\in\mathbb{R}^{n}$,
    $\sup_{\varphi\in\mathcal{V}}\left|\sum_{i=1}^{n}a_{i,n}(\varphi)x_{i:n}-\int_{0}^{1}\varphi(\alpha)q_{n}(\alpha)d\alpha\right|\leqslant\frac{L}{n}\int_{0}^{1}|q_{n}(\alpha)|d\alpha.$
Proof.
(i) For $t\in((i-1)/n,i/n]$, we have $\varphi_{n}(t)=n\int_{(i-1)/n}^{i/n}\varphi(s)ds$, the average of $\varphi$ on that interval. For any $s\in((i-1)/n,i/n]$ and Lipschitz $\varphi$,
$|\varphi(s)-\varphi(t)|\leqslant L|s-t|\leqslant L/n.$
Thus $|\varphi_{n}(t)-\varphi(t)|\leqslant L/n$, uniformly in $t$ and $\varphi\in\mathcal{V}$.
(ii) By Lemma 6.5, $\sum_{i=1}^{n}a_{i,n}(\varphi)x_{i:n}=\int_{0}^{1}\varphi_{n}(\alpha)q_{n}(\alpha)d\alpha$. Thus
$\left|\sum_{i=1}^{n}a_{i,n}(\varphi)x_{i:n}-\int_{0}^{1}\varphi(\alpha)q_{n}(\alpha)d\alpha\right|$ $=\left|\int_{0}^{1}(\varphi_{n}-\varphi)q_{n}\,d\alpha\right|$
$\leqslant\|\varphi_{n}-\varphi\|_{\infty}\int_{0}^{1}|q_{n}|$
$\leqslant\frac{L}{n}\int_{0}^{1}|q_{n}|.\qed$
Theorem 7.6 (Uniform spectral plug-in consistency).
Let $\mathcal{V}$ be a Lipschitz spectral class with constants $(C,L)$. Let $X\in L^{1}$ and $(X_{i})$ be i.i.d. with law $X$. Then
(23) $\sup_{\varphi\in\mathcal{V}}\left|\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right|\xrightarrow[n\to\infty]{\mathrm{a.s.}}0.$
(The uniformity is over spectra $\varphi\in\mathcal{V}$ for a fixed underlying law of $X$.)
Proof.
Let $q_{n}(\alpha):=X_{\lceil n\alpha\rceil:n}$ be the empirical quantile function and let $q_{X}$ be the population lower quantile. For each $\varphi\in\mathcal{V}$, write $\varphi_{n}$ for the step function associated with the canonical weights (Definitions 6.2 and 6.4). By Lemma 6.5,
$\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})=-\int_{0}^{1}\varphi_{n}(\alpha)\,q_{n}(\alpha)\,d\alpha,\qquad\rho_{\varphi}(X)=-\int_{0}^{1}\varphi(\alpha)\,q_{X}(\alpha)\,d\alpha.$
Hence,
$\sup_{\varphi\in\mathcal{V}}\left|\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right|$ $\leqslant\sup_{\varphi\in\mathcal{V}}\left|\int_{0}^{1}(\varphi_{n}-\varphi)\,q_{n}\,d\alpha\right|+\sup_{\varphi\in\mathcal{V}}\left|\int_{0}^{1}\varphi\,(q_{n}-q_{X})\,d\alpha\right|.$
For the first term, Lemma 7.5(i) yields
$\sup_{\varphi\in\mathcal{V}}\left|\int_{0}^{1}(\varphi_{n}-\varphi)\,q_{n}\,d\alpha\right|\leqslant\frac{L}{n}\int_{0}^{1}|q_{n}(\alpha)|\,d\alpha=\frac{L}{n}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}|.$
By the strong law, $\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\to\mathsf{E}|X|$ almost surely, so this term tends to 0 almost surely.
For the second term, boundedness of the class gives
$\sup_{\varphi\in\mathcal{V}}\left|\int_{0}^{1}\varphi\,(q_{n}-q_{X})\,d\alpha\right|\leqslant\sup_{\varphi\in\mathcal{V}}\|\varphi\|_{\infty}\int_{0}^{1}|q_{n}-q_{X}|\leqslant C\int_{0}^{1}|q_{n}-q_{X}|.$
It remains to show $\int_{0}^{1}|q_{n}-q_{X}|\,d\alpha\to 0$ almost surely when $X\in L^{1}$. By Glivenko–Cantelli, $q_{n}(\alpha)\to q_{X}(\alpha)$ at continuity points of $q_{X}$, hence almost everywhere. Fix $M>0$ and define truncations $q_{n}^{(M)}$ and $q_{X}^{(M)}$ by clamping to [M,M][-M,M]. Then $q_{n}^{(M)}\to q_{X}^{(M)}$ almost everywhere and $|q_{n}^{(M)}-q_{X}^{(M)}|\leqslant 2M$, so dominated convergence gives $\int_{0}^{1}|q_{n}^{(M)}-q_{X}^{(M)}|\,d\alpha\to 0$ almost surely for fixed $M$. Moreover,
$\int_{0}^{1}|q_{n}-q_{X}|\leqslant\int_{0}^{1}|q_{n}-q_{n}^{(M)}|+\int_{0}^{1}|q_{n}^{(M)}-q_{X}^{(M)}|+\int_{0}^{1}|q_{X}^{(M)}-q_{X}|.$
The middle term vanishes as $n\to\infty$. Since $X\in L^{1}$, we have $q_{X}\in L^{1}(0,1)$, so $\int_{0}^{1}|q_{X}^{(M)}-q_{X}|\to 0$ as $M\to\infty$. Finally,
$\int_{0}^{1}|q_{n}(\alpha)|\,\mathbf{1}_{\{|q_{n}(\alpha)|>M\}}\,d\alpha=\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\,\mathbf{1}_{\{|X_{i}|>M\}}\longrightarrow\mathsf{E}[|X|\,\mathbf{1}_{\{|X|>M\}}]$
almost surely by the strong law, and the limit tends to 0 as $M\to\infty$. Thus $\int_{0}^{1}|q_{n}-q_{X}|\to 0$ almost surely, and the second term tends to 0 almost surely as well.
Combining the two bounds proves (23). ∎
Remark 7.7.
Theorem 7.6 requires spectra in the Lipschitz class $\mathcal{V}$. This covers distortion risk measures and spectral measures with smooth weights, but does not cover the Expected Shortfall spectrum $\varphi_{\alpha}(u)=\alpha^{-1}\mathbf{1}_{(0,\alpha]}(u)$, which is bounded but not Lipschitz (it has a jump discontinuity at $u=\alpha$).
For ES specifically, one can prove consistency by a direct argument exploiting the explicit formula $\operatorname{ES}_{\alpha}=-\frac{1}{\alpha}\int_{0}^{\alpha}q(u)\,du$ and Glivenko–Cantelli, but the uniform rate over $\alpha\in[\delta,1]$ requires additional care; see [7] for sharp uniform results on ES estimation.
Corollary 7.8.
Under the conditions of Theorem 7.6, assume additionally that $\mathsf{E}[|X|^{2+\eta}]<\infty$ for some $\eta>0$. Then
$\sup_{\varphi\in\mathcal{V}}\left|\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right|=O_{\mathsf{P}}\!\left(\frac{1}{\sqrt{n}}\right).$
If, in addition, $X$ is almost surely bounded and the distribution function $F$ is continuous, then
$\sup_{\varphi\in\mathcal{V}}\left|\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right|=O\!\left(\sqrt{\frac{\log\log n}{n}}\right)\quad\text{almost surely.}$
Proof.
From the proof of Theorem 7.6 we have $\sup_{\varphi\in\mathcal{V}}|\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X)|\leqslant A_{n}+B_{n}$ with
$A_{n}\leqslant\frac{L}{n}\int_{0}^{1}|q_{n}(\alpha)|\,d\alpha=\frac{L}{n}\cdot\frac{1}{n}\sum_{i=1}^{n}|X_{i}|\quad\text{and}\quad B_{n}\leqslant C\int_{0}^{1}|q_{n}(\alpha)-q_{X}(\alpha)|\,d\alpha=C\,W_{1}(\hat{\mu}_{n},\mu).$
(The quantile formula $W_{1}(\hat{\mu}_{n},\mu)=\int_{0}^{1}|q_{n}-q_{X}|$ holds whenever both measures have finite first moments—guaranteed here by $X\in L^{1}$ from Theorem 7.6; see, e.g., [20, Theorem 2.18].) Thus $A_{n}=O(n^{-1})$ almost surely by the strong law.
For the stochastic term, the one-dimensional identity $W_{1}(\hat{\mu}_{n},\mu)=\int_{\mathbb{R}}|F_{n}(x)-F(x)|\,dx$ (again [20]) converts the quantile integral to a CDF integral. For each fixed $x$, conditional on $F(x)$ we have $F_{n}(x)=\frac{1}{n}\mathrm{Bin}(n,F(x))$, hence $\mathsf{E}|F_{n}(x)-F(x)|\leqslant\sqrt{\operatorname{\mathsf{Var}}(F_{n}(x))}=\sqrt{F(x)(1-F(x))/n}$. By the Fubini theorem and Jensen’s inequality,
$\mathsf{E}\,W_{1}(\hat{\mu}_{n},\mu)\leqslant\frac{1}{\sqrt{n}}\int_{\mathbb{R}}\sqrt{F(x)(1-F(x))}\,dx.$
The integral is finite under $\mathsf{E}|X|^{2+\eta}<\infty$. Indeed, for $x\geqslant 1$,
$1-F(x)=\mathsf{P}(X>x)\leqslant\frac{\mathsf{E}|X|^{2+\eta}}{x^{2+\eta}}$
by Markov’s inequality, and hence
$\sqrt{F(x)(1-F(x))}\leqslant\sqrt{1-F(x)}\leqslant\sqrt{\mathsf{E}|X|^{2+\eta}}\,x^{-(1+\eta/2)}.$
Since $\eta>0$, the function $x\mapsto x^{-(1+\eta/2)}$ is integrable on $1,\infty)$. A symmetric argument for $x\leqslant-1$ (using $\mathsf{P}(X\leqslant x)=\mathsf{P}(-X\geqslant-x)$) yields integrability on $(-\infty,-1]$, and integrability on [1,1][-1,1] is trivial because $\sqrt{F(1-F)}\leqslant 1/2$. Therefore $\int_{\mathbb{R}}\sqrt{F(x)(1-F(x))}\,dx<\infty$.
Hence $\mathsf{E}W_{1}(\hat{\mu}_{n},\mu)=O(n^{-1/2})$, and Markov’s inequality yields $W_{1}(\hat{\mu}_{n},\mu)=O_{\mathsf{P}}(n^{-1/2})$. (This rate is standard in the 1D empirical Wasserstein literature; see, e.g., [5] for optimal moment conditions.) Hence $B_{n}=O_{\mathsf{P}}(n^{-1/2})$, and the first claim follows.
If $X$ is bounded and $F$ is continuous, then the probability integral transform and the law of the iterated logarithm for the uniform empirical process give $\sup_{x}|F_{n}(x)-F(x)|=O(\sqrt{\log\log n/n})$ almost surely; see, e.g., [19]. Since $X$ is bounded, $W_{1}(\hat{\mu}_{n},\mu)=\int_{\mathbb{R}}|F_{n}-F|\,dx$ is bounded by $\mathrm{diam}(\mathrm{supp}(X))\sup_{x}|F_{n}(x)-F(x)|$, giving the stated almost sure rate. ∎
Remark 7.9.
From the hyperfinite viewpoint, Theorem 7.6 reflects the fact that one can work internally with the entire class $\mathcal{V}$ at once. The Lipschitz condition provides a deterministic envelope bounding the discretisation error uniformly, whilst the Loeb-measure quantile convergence handles the stochastic approximation uniformly over all bounded $\varphi$.
8. Kusuoka-type plug-in consistency
We now extend the consistency theory to general law-invariant coherent risk measures via the Kusuoka representation.
8.1. Setup and assumptions
Let $\rho:L^{\infty}\to\mathbb{R}$ be a law-invariant CRM on an atomless space with Kusuoka representation (3):
$\rho(X)=\sup_{\nu\in\mathcal{M}}\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha).$
We construct finite-sample estimators by discretising the Kusuoka integral. Let $(\alpha_{i,n})_{i=0}^{n}$ be a grid with $0=\alpha_{0,n}<\alpha_{1,n}<\cdots<\alpha_{n,n}=1$. For simplicity, take the uniform grid $\alpha_{i,n}=i/n$.
Suppose we have, for each $\alpha\in(0,1]$, an estimator $\widehat{\operatorname{ES}}_{\alpha,n}:\mathbb{R}^{n}\to\mathbb{R}$ for $\operatorname{ES}_{\alpha}$. The Kusuoka-type plug-in estimator is
(24) $\hat{\rho}_{n}(x):=\sup_{\nu\in\mathcal{M}}\sum_{i=1}^{n}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(x)\,\nu((\alpha_{i-1,n},\alpha_{i,n}]).$
8.2. Consistency theorem
Theorem 8.1 (Kusuoka plug-in consistency).
Let $\rho$ be a law-invariant CRM with Kusuoka representation (3). Assume:
  1. (K1)
    (Tightness) For every $\varepsilon>0$, there exists $\delta\in(0,1)$ such that $\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\varepsilon$.
  2. (K2)
    (Uniform ES estimation) For every $\delta\in(0,1)$,
    $\sup_{\alpha\in[\delta,1]}\left|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha}(X)\right|\xrightarrow[n\to\infty]{\mathrm{a.s.}}0$
    for any bounded $X$ and i.i.d. sample $(X_{i})$ with law $X$.
  3. (K3)
    (Uniform envelope) There exists $C<\infty$ such that $|\widehat{\operatorname{ES}}_{\alpha,n}(x)|\leqslant C\|x\|_{\infty}$ for all $\alpha,n,x$.
  4. (K4)
    (Grid refinement) The mesh $\max_{i}(\alpha_{i,n}-\alpha_{i-1,n})\to 0$ as $n\to\infty$.
Then for any bounded $X$ and i.i.d. sample $(X_{i})$,
(25) $\hat{\rho}_{n}(X_{1},\ldots,X_{n})\xrightarrow[n\to\infty]{\mathrm{a.s.}}\rho(X).$
Assumption (K1) excludes risk measures that put mass arbitrarily close to $\alpha=0$, which is necessary because $\operatorname{ES}_{\alpha}(X)\to$-$\infty$ as $\alpha\downarrow 0$ for unbounded-below $X$, and discretisation near 0 is unstable. A typical choice for $\widehat{\operatorname{ES}}_{\alpha,n}$ is the discrete ES $\mathrm{dES}_{\lceil n\alpha\rceil/n}$ from (13), for which (K2)–(K3) hold; see Theorem 7.2.
Proof.
Fix a bounded random variable $X$ and an i.i.d. sample $(X_{i})_{i\geqslant 1}$ with law $X$. For $\nu\in\mathcal{M}$ set
$I(\nu):=\int_{(0,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha),\qquad\hat{I}_{n}(\nu):=\sum_{i=1}^{n}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}]).$
Then $\rho(X)=\sup_{\nu\in\mathcal{M}}I(\nu)$ and $\hat{\rho}_{n}(X_{1},\ldots,X_{n})=\sup_{\nu\in\mathcal{M}}\hat{I}_{n}(\nu)$. Since $|\sup_{\nu}a(\nu)-\sup_{\nu}b(\nu)|\leqslant\sup_{\nu}|a(\nu)-b(\nu)|$, it suffices to prove that
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\longrightarrow 0\qquad\text{almost surely.}$
Fix $\varepsilon>0$ and choose $\delta\in(0,1)$ such that $\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\varepsilon$ (assumption (K1)). For each $\nu\in\mathcal{M}$, split
$I(\nu)=\int_{(0,\delta]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)+\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)=:I_{0,\delta}(\nu)+I_{\delta,1}(\nu),$
and split $\hat{I}_{n}(\nu)$ analogously into the terms with $\alpha_{i,n}\leqslant\delta$ and $\alpha_{i,n}>\delta$:
$\hat{I}_{n}(\nu)=\sum_{\alpha_{i,n}\leqslant\delta}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}])\\ +\sum_{\alpha_{i,n}>\delta}\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})\,\nu((\alpha_{i-1,n},\alpha_{i,n}])=:\hat{I}_{0,\delta,n}(\nu)+\hat{I}_{\delta,1,n}(\nu).$
Because $X$ is bounded we have $|\operatorname{ES}_{\alpha}(X)|\leqslant\|X\|_{\infty}$ for all $\alpha\in(0,1]$. Therefore,
$\sup_{\nu\in\mathcal{M}}|I_{0,\delta}(\nu)|\leqslant\|X\|_{\infty}\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant\|X\|_{\infty}\,\varepsilon.$
Moreover, since $X$ is bounded, $\|(X_{1},\ldots,X_{n})\|_{\infty}\leqslant\|X\|_{\infty}$ almost surely, and assumption (K3) yields $|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})|\leqslant C\|X\|_{\infty}$ for all $\alpha,n$ on an almost sure event. Hence,
$\sup_{\nu\in\mathcal{M}}|\hat{I}_{0,\delta,n}(\nu)|\leqslant C\|X\|_{\infty}\sup_{\nu\in\mathcal{M}}\nu((0,\delta])\leqslant C\|X\|_{\infty}\,\varepsilon\qquad\text{almost surely.}$
Combining these two bounds gives
$\sup_{\nu\in\mathcal{M}}\bigl|\hat{I}_{0,\delta,n}(\nu)-I_{0,\delta}(\nu)\bigr|\leqslant(1+C)\|X\|_{\infty}\,\varepsilon\qquad\text{almost surely.}$
It remains to control the contribution from $(\delta,1]$. Write
$\hat{I}_{\delta,1,n}(\nu)-I_{\delta,1}(\nu)$ $=\sum_{\alpha_{i,n}>\delta}\Big(\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha_{i,n}}(X)\Big)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])$
$\quad+\sum_{\alpha_{i,n}>\delta}\operatorname{ES}_{\alpha_{i,n}}(X)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])-\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha).$
For the first term we simply use $\nu((\delta,1])\leqslant 1$ and assumption (K2):
$\sup_{\nu\in\mathcal{M}}\left|\sum_{\alpha_{i,n}>\delta}\Big(\widehat{\operatorname{ES}}_{\alpha_{i,n},n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha_{i,n}}(X)\Big)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])\right|\\ \leqslant\sup_{\alpha\in[\delta,1]}\left|\widehat{\operatorname{ES}}_{\alpha,n}(X_{1},\ldots,X_{n})-\operatorname{ES}_{\alpha}(X)\right|\longrightarrow 0$
almost surely.
For the second term, define $g(\alpha):=\operatorname{ES}_{\alpha}(X)$ on $\delta,1]$. Since $X$ is bounded, $q_{X}$ is bounded, so $\alpha\mapsto\int_{0}^{\alpha}q_{X}(u)\,du$ is continuous and therefore $g(\alpha)=-(1/\alpha)\int_{0}^{\alpha}q_{X}(u)\,du$ is continuous on $\delta,1]$. By compactness, $g$ is uniformly continuous on $\delta,1]$. Let $\omega_{g}(h)$ denote its modulus of continuity:
$\omega_{g}(h):=\sup\{|g(\alpha)-g(\beta)|:\ \alpha,\beta\in[\delta,1],\ |\alpha-\beta|\leqslant h\}.$
Let $\Delta_{n}:=\max_{i}(\alpha_{i,n}-\alpha_{i-1,n})$, which tends to 0 by (K4). Then, for every $\nu\in\mathcal{M}$,
$\left|\sum_{\alpha_{i,n}>\delta}\operatorname{ES}_{\alpha_{i,n}}(X)\,\nu((\alpha_{i-1,n},\alpha_{i,n}])-\int_{(\delta,1]}\operatorname{ES}_{\alpha}(X)\,\nu(d\alpha)\right|$
$\qquad=\left|\sum_{\alpha_{i,n}>\delta}\int_{(\alpha_{i-1,n},\alpha_{i,n}]}\bigl(g(\alpha_{i,n})-g(\alpha)\bigr)\,\nu(d\alpha)\right|$
$\qquad\leqslant\sum_{\alpha_{i,n}>\delta}\int_{(\alpha_{i-1,n},\alpha_{i,n}]}\omega_{g}(\Delta_{n})\,\nu(d\alpha)\leqslant\omega_{g}(\Delta_{n}).$
This bound is uniform in $\nu$, and $\omega_{g}(\Delta_{n})\to 0$ as $n\to\infty$.
Putting everything together, we obtain on an almost sure event:
$\limsup_{n\to\infty}\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\leqslant(1+C)\|X\|_{\infty}\,\varepsilon.$
Since $\varepsilon>0$ is arbitrary, $\sup_{\nu\in\mathcal{M}}|\hat{I}_{n}(\nu)-I(\nu)|\to 0$ almost surely, we conclude convergence $\hat{\rho}_{n}(X_{1},\ldots,X_{n})\to\rho(X)$ a.s. ∎
9. Hyperfinite bootstrap validity
We establish bootstrap validity by exploiting the internal nature of the resampling procedure. In the hyperfinite framework, the bootstrap distribution is simply the internal distribution of the statistic conditional on the hyperfinite sample.
9.1. Internal bootstrap construction
Let $(X_{1},\dots,X_{N})$ be a hyperfinite sample with infinite $N\in{}^{*}\mathbb{N}$. The empirical measure is $\hat{\mathsf{P}}_{N}=\frac{1}{N}\sum_{k=1}^{N}\delta_{X_{k}}$.
Introduce an internal i.i.d. sequence $(U_{1},\dots,U_{N})$, independent of $(X_{1},\dots,X_{N})$, with each $U_{i}$ uniformly distributed on $I_{N}=\{1,\dots,N\}$ (counting law $\mu_{N}$). Define the hyperfinite bootstrap sample by
$X_{i}^{*}:=X_{U_{i}},\qquad i=1,\dots,N.$
Then $(X_{1}^{*},\dots,X_{N}^{*})$ is internal and i.i.d. with law $\hat{\mathsf{P}}_{N}$, and conditional-on-sample statements are interpreted as statements under the bootstrap-index randomness $(U_{i})$ with $(X_{i})$ held fixed. This construction makes the internal conditional law explicit: it is simply the pushforward of the internal product counting law of $(U_{i})$.
By the star-suppression convention (Notation 3.2), the mapping $(n,x)\mapsto\hat{\rho}_{n,\varphi}(x)$ extends internally to hypernatural $N$, giving the internal statistic $\hat{\rho}_{N,\varphi}$. Let $\hat{\rho}_{N}:=\hat{\rho}_{N,\varphi}(X_{1},\dots,X_{N})$ be the estimator on the original sample, and $\hat{\rho}_{N}^{*}:=\hat{\rho}_{N,\varphi}(X_{1}^{*},\dots,X_{N}^{*})$ be the estimator on the bootstrap sample.
We measure closeness of laws using the Kolmogorov distance. For standard probability measures,
$d_{K}(\mathcal{L}(Y),\mathcal{L}(Z)):=\sup_{t\in\mathbb{R}}\big|\mathsf{P}(Y\leqslant t)-\mathsf{P}(Z\leqslant t)\big|.$
In the hyperfinite setting, we must be careful about internality. For $m\in{}^{*}\mathbb{N}$, define the internal truncated Kolmogorov distance
$d_{K,m}(\mathcal{L}(Y),\mathcal{L}(Z)):=\max_{t\in\{-m,-m+1/m,\dots,m\}}\big|{}^{*}\mathsf{P}(Y\leqslant t)-{}^{*}\mathsf{P}(Z\leqslant t)\big|.$
This is internal (a maximum over an internal hyperfinite grid).
Lemma 9.1 (Internal Kolmogorov distance approximation).
Let $Y$ be a finite ${}^{*}\mathbb{R}$-valued random variable on $({}^{*}\Omega,L({}^{*}\mathsf{P}))$, and let $\mu$ be a standard probability measure on $\mathbb{R}$ with continuous CDF $G$. For $m\in{}^{*}\mathbb{N}$ define
$d_{K,m}(\mathcal{L}^{*}(Y),\mu):=\max_{t\in\{-m,-m+1/m,\dots,m\}}\left|\mathcal{L}^{*}(Y)((-\infty,t])-G(t)\right|.$
  1. (i)
    If there exists an infinite $m\in{}^{*}\mathbb{N}$ such that $d_{K,m}(\mathcal{L}^{*}(Y),\mu)\approx 0$, then
    $d_{K}(\mathcal{L}(\operatorname{st}(Y)),\mu)=0.$
  2. (ii)
    The event $\{d_{K,m}(\mathcal{L}^{*}(Y\mid X),\mu)>\varepsilon\}$ is internal for each internal $m\in{}^{*}\mathbb{N}$ and each standard $\varepsilon>0$, hence Loeb measurable.
Proof.
(i) Fix an infinite $m\in{}^{*}\mathbb{N}$ with $d_{K,m}(\mathcal{L}^{*}(Y),\mu)\approx 0$ and write $\Delta:=1/m$. Let $t\in\mathbb{R}$ be standard and let $\eta>0$ be standard. Since $m$ is infinite, we have $t,t+\eta\in(-m,m)$. Choose gridpoints $s,s^{\prime}\in\{-m,-m+1/m,\dots,m\}$ such that
$s\leqslant t<s+\Delta,\qquad s^{\prime}\leqslant t+\eta<s^{\prime}+\Delta.$
Then
$\{Y\leqslant s\}\subseteq\{\operatorname{st}(Y)\leqslant t\}\subseteq\{Y\leqslant t+\eta\}\subseteq\{Y\leqslant s^{\prime}+\Delta\}.$
Taking Loeb probabilities and using $L({}^{*}\mathsf{P})(A)=\operatorname{st}({}^{*}\mathsf{P}(A))$ for internal $A$ gives
$\operatorname{st}({}^{*}\mathsf{P}(Y\leqslant s))\leqslant L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t)\leqslant\operatorname{st}({}^{*}\mathsf{P}(Y\leqslant s^{\prime}+\Delta)).$
By the definition of $d_{K,m}$,
$\big|{}^{*}\mathsf{P}(Y\leqslant s)-G(s)\big|\leqslant d_{K,m}(\mathcal{L}^{*}(Y),\mu)\approx 0,$
and
$\big|{}^{*}\mathsf{P}(Y\leqslant s^{\prime}+\Delta)-G(s^{\prime}+\Delta)\big|\leqslant d_{K,m}(\mathcal{L}^{*}(Y),\mu)\approx 0.$
Hence
$\operatorname{st}(G(s))\leqslant L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t)\leqslant\operatorname{st}(G(s^{\prime}+\Delta)).$
Since $s\approx t$ and $s^{\prime}+\Delta\approx t+\eta$ and $G$ is continuous, we have $\operatorname{st}(G(s))=G(t)$ and $\operatorname{st}(G(s^{\prime}+\Delta))=G(t+\eta)$. Therefore
$G(t)\leqslant L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t)\leqslant G(t+\eta).$
Letting $\eta\downarrow 0$ and using continuity of $G$ yields $L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t)=G(t)$ for all $t\in\mathbb{R}$, i.e. $\operatorname{st}(Y)\sim\mu$.
(ii) For fixed internal $m$, the grid $\{-m,-m+1/m,\dots,m\}$ is hyperfinite and $d_{K,m}$ is a hyperfinite maximum of internal conditional probabilities and standard constants. Thus $\{d_{K,m}(\mathcal{L}^{*}(Y\mid X),\mu)>\varepsilon\}$ is internal for standard $\varepsilon>0$. ∎
9.2. Near-standardness of conditional laws
Bootstrap validity requires that the conditional law of $\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})$ is infinitely close to the limiting law of $\sqrt{N}(\hat{\rho}_{N}-\rho(X))$.
Assumption 9.2 (Bootstrap regularity).
The random variable $X$ satisfies:
  1. (B1)
    $F_{X}$ is continuous;
  2. (B2)
    $X$ has a density $f_{X}$ that is bounded and bounded away from 0 on the interior quantile range $\{q_{X}(\alpha):\alpha\in[\delta,1-\delta]\}$ for each fixed $\delta\in(0,1/2)$.
Equivalently, (B2) says $f_{X}$ is continuous and strictly positive on every compact subinterval of the interior of the support of $X$. Under (B1)–(B2), the quantile map $F\mapsto q_{F}$ is Hadamard differentiable at $F_{X}$, and the spectral functional $T(F)=-\int_{0}^{1}q_{F}(\alpha)\varphi(\alpha)\,d\alpha$ inherits Hadamard differentiability; see [18, Lemma 21.3, Thm. 21.5].
Lemma 9.3 (Hadamard derivative of the spectral functional).
Under Assumption 9.2, the spectral functional $T(F)=-\int_{0}^{1}q_{F}(\alpha)\varphi(\alpha)\,d\alpha$ is Hadamard differentiable at $F_{X}$ tangentially to $C(\mathbb{R})$, with derivative
$T^{\prime}_{F_{X}}(h)=\int_{0}^{1}\frac{h(q_{X}(\alpha))}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha.$
Proof.
The quantile functional $Q:F\mapsto q_{F}(\alpha)$ has Hadamard derivative $Q^{\prime}_{F_{X}}(h)(\alpha)=-h(q_{X}(\alpha))/f_{X}(q_{X}(\alpha))$ at continuity points $\alpha$ of $q_{X}$; see [18, Lemma 21.1]. Since
$T(F)=-\int_{0}^{1}Q(F)(\alpha)\varphi(\alpha)\,d\alpha,$
the chain rule gives
$T^{\prime}_{F_{X}}(h)=-\int_{0}^{1}Q^{\prime}_{F_{X}}(h)(\alpha)\,\varphi(\alpha)\,d\alpha=\int_{0}^{1}\frac{h(q_{X}(\alpha))}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha.\qed$
Theorem 9.4 (Hyperfinite Bootstrap Consistency).
Let $X\in L^{2}$ satisfy Assumption 9.2, and let $\varphi$ be Lipschitz. Assume moreover that the hypotheses of Theorem 10.5 hold, so that the asymptotic variance $\sigma_{\varphi}^{2}$ in (28) is finite. Let $\mathcal{L}^{*}(\cdot|X)$ denote the internal conditional law given the sample (i.e., the law under the bootstrap-index randomness $(U_{i})$ with $(X_{i})$ fixed). Then, for $L({}^{*}\mathsf{P})$-almost all sample paths $\omega$,
$d_{K}\left(\mathcal{L}^{*}\left(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\;\Big|\;X(\omega)\right),\ N(0,\sigma_{\varphi}^{2})\right)\approx 0,$
where $\sigma_{\varphi}^{2}$ is the asymptotic variance in Theorem 10.5.
Proof.
The proof combines the standard functional delta method for the bootstrap as explained by [18, Thm. 23.9] with NSA reasoning, using the internal Kolmogorov distance $d_{K,m}$ from Lemma 9.1 to ensure measurability.
Near-standardness of the empirical measure. Fix a sample path $\omega$ such that the empirical CDF $\hat{F}_{N}$ satisfies $\|\hat{F}_{N}-F_{X}\|_{\infty}\approx 0$. By the hyperfinite Glivenko–Cantelli theorem (Theorem 3.20), this holds for $L({}^{*}\mathsf{P})$-almost all $\omega$.
Standard bootstrap delta method. The standard bootstrap delta method [18, Thm. 23.9] asserts: under Assumption 9.2, the conditional distribution of $\sqrt{n}(T(\hat{F}_{n}^{*})-T(\hat{F}_{n}))$ given the sample converges to $\mathcal{L}(T^{\prime}_{F_{X}}(\mathbb{G}))$ in probability in the bounded-Lipschitz metric, where $\mathbb{G}$ is an $F_{X}$-Brownian bridge. Since the limit distribution $N(0,\sigma_{\varphi}^{2})$ has a continuous CDF, this implies convergence in Kolmogorov distance [18, Lemma 21.2]: for every $\varepsilon,\delta>0$ and $m\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
$\mathsf{P}\Big(d_{K,m}\big(\mathcal{L}(\sqrt{n}(T(\hat{F}_{n}^{*})-T(\hat{F}_{n}))\mid X),\,N(0,\sigma_{\varphi}^{2})\big)>\varepsilon\Big)<\delta.$
NSA repackaging via internal distance. Fix standard $\varepsilon,\delta>0$ and standard $m\in\mathbb{N}$. The event $E_{m,\varepsilon}:=\{d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))>\varepsilon\}$ is internal by Lemma 9.1(ii), hence Loeb measurable.
The sample size $N$ exceeds any standard threshold $n_{0}$ (since $N$ is infinite), so $^{*}\mathsf{P}(E_{m,\varepsilon})<\delta$. Taking Loeb measures and using that $\delta>0$ is arbitrary, $L({}^{*}\mathsf{P})(E_{m,\varepsilon})=0$ for each standard $\varepsilon>0$ and $m\in\mathbb{N}$.
Intersecting over countably many $\varepsilon=1/k$ and $m\in\mathbb{N}$, we obtain a Loeb-probability-one set $G$ on which $d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))<1/r$ for all standard $m,r\in\mathbb{N}$. To obtain an infinite $M$ with $d_{K,M}\approx 0$, we use saturation: for each standard $r\in\mathbb{N}$, define the internal set
$A_{r}:=\{m\in{}^{*}\mathbb{N}:d_{K,m}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))<1/r\}.$
On the set $G$, each $A_{r}$ contains all standard $m$. The family $\{A_{r}\}_{r\in\mathbb{N}}\cup\{B_{s}\}_{s\in\mathbb{N}}$ (where $B_{s}:=\{m\in{}^{*}\mathbb{N}:m>s\}$) has the finite intersection property. By countable saturation, there exists $M\in\bigcap_{r\in\mathbb{N}}A_{r}\cap\bigcap_{s\in\mathbb{N}}B_{s}$, which is infinite (since $M>s$ for all standard $s$) and satisfies $d_{K,M}<1/r$ for all standard $r$, hence $d_{K,M}\approx 0$. By Lemma 9.1(i), the standard-part distribution equals $N(0,\sigma_{\varphi}^{2})$.
Identification of the asymptotic variance. The random variable $T^{\prime}_{F_{X}}(\mathbb{G})$ equals
$\int_{0}^{1}\frac{\mathbb{G}(q_{X}(\alpha))}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha=\int_{0}^{1}\frac{B(\alpha)}{f_{X}(q_{X}(\alpha))}\,\varphi(\alpha)\,d\alpha,$
where $B(\alpha):=\mathbb{G}(q_{X}(\alpha))$ is a standard Brownian bridge on $[0,1]$ (since $\mathbb{G}(x)$ at $x=q_{X}(\alpha)$ with $F_{X}(q_{X}(\alpha))=\alpha$ gives $\mathrm{Cov}(B(\alpha),B(\beta))=\min(\alpha,\beta)-\alpha\beta$).
This integral is Gaussian with mean zero. Under the stronger smoothness hypotheses of Theorem 10.5 (in particular, $q_{X}\in C^{1}$ with $q_{X}^{\prime}=1/f_{X}(q_{X})$), the variance equals
$\sigma_{\varphi}^{2}=\int_{0}^{1}\int_{0}^{1}(\min(\alpha,\beta)-\alpha\beta)\,\frac{\varphi(\alpha)\varphi(\beta)}{f_{X}(q_{X}(\alpha))f_{X}(q_{X}(\beta))}\,d\alpha\,d\beta.$
Since the limit $N(0,\sigma_{\varphi}^{2})$ has a continuous CDF, weak convergence to it implies Kolmogorov distance convergence (see, e.g., [18, Lemma 21.2]). Thus $d_{K}(\mathcal{L}^{*}(\sqrt{N}(\hat{\rho}_{N}^{*}-\hat{\rho}_{N})\mid X),N(0,\sigma_{\varphi}^{2}))\approx 0$. ∎
Remark 9.5 (The power of the hyperfinite viewpoint).
The advantage of this approach is that we do not need to check “sequences” of conditional measures converging in a double-limit sense. We fix a single infinite $N$ and a typical path; the bootstrap distribution is then a fixed internal distribution, which we show is near-standard to the Gaussian. The transfer principle converts the standard Hadamard differentiability into an internal statement, and near-standardness of the empirical measure (from Theorem 3.20) provides the bridge between the internal and standard worlds.
10. Asymptotic normality via the hyperfinite CLT
We derive the asymptotic distribution of spectral plug-in estimators using the hyperfinite central limit theorem.
10.1. Hyperfinite central limit theorem
Theorem 10.1 (Hyperfinite CLT).
Let $(X_{i})_{i\in\mathbb{N}}$ be i.i.d. on $(\Omega,\mathscr{G},\mathsf{P})$ with $\mathsf{E}[X_{1}]=0$ and $\operatorname{\mathsf{Var}}(X_{1})=\sigma^{2}<\infty$. Let $N\in{}^{*}\mathbb{N}$ be infinite and consider the hyperfinite sum $S_{N}:=\frac{1}{\sqrt{N}}\sum_{k=1}^{N}X_{k}$ on $({}^{*}\Omega,L({}^{*}\mathscr{G}),L({}^{*}\mathsf{P}))$. Under the finite variance assumption, $S_{N}$ is finite $L({}^{*}\mathsf{P})$-almost surely. Indeed, by Chebyshev’s inequality (which transfers), for any standard $M>0$,
$^{*}\mathsf{P}(|S_{N}|>M)\leqslant\frac{{}^{*}\mathsf{E}[S_{N}^{2}]}{M^{2}}=\frac{\sigma^{2}}{M^{2}}.$
Taking standard parts, $L({}^{*}\mathsf{P})(|S_{N}|>M)\leqslant\sigma^{2}/M^{2}$. Letting $M\to\infty$,
$L({}^{*}\mathsf{P})(|S_{N}|\text{ is infinite})=0.$
Thus $\operatorname{st}(S_{N})$ is well-defined $L({}^{*}\mathsf{P})$-almost surely.
Then for every $t\in\mathbb{R}$,
$L({}^{*}\mathsf{P})\bigl(\operatorname{st}(S_{N})\leqslant t\bigr)=\Phi(t/\sigma),$
where $\Phi$ is the standard normal distribution function.
Remark 10.2.
The hyperfinite CLT is a standard result in non-standard probability; see, e.g., [3, Ch. 5] or [4] for systematic treatments.
Proof.
By the classical CLT, $S_{n}:=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}X_{k}\xrightarrow{d}N(0,\sigma^{2})$. For any $t\in\mathbb{R}$ and $\varepsilon>0$, there exists $n_{0}$ such that for all $n\geqslant n_{0}$,
$|\mathsf{P}(S_{n}\leqslant t)-\Phi(t/\sigma)|<\varepsilon.$
By transfer, this holds for $N$, hence $|{}^{*}\mathsf{P}(S_{N}\leqslant t)-\Phi(t/\sigma)|<\varepsilon$. Since $\varepsilon$ is arbitrary, $^{*}\mathsf{P}(S_{N}\leqslant t)\approx\Phi(t/\sigma)$.
To pass from $^{*}\mathsf{P}(S_{N}\leqslant t)$ to the Loeb event $\{\operatorname{st}(S_{N})\leqslant t\}$, note that
$\{\operatorname{st}(S_{N})\leqslant t\}=\bigcap_{m\in\mathbb{N}}\{S_{N}\leqslant t+1/m\},$
hence by continuity from above of $L({}^{*}\mathsf{P})$,
$L({}^{*}\mathsf{P})(\operatorname{st}(S_{N})\leqslant t)=\lim_{m\to\infty}\operatorname{st}\bigl({}^{*}\mathsf{P}(S_{N}\leqslant t+1/m)\bigr).$
Since $^{*}\mathsf{P}(S_{N}\leqslant t+1/m)\approx\Phi((t+1/m)/\sigma)$ for each fixed $m$ and $\Phi$ is continuous, the limit equals $\Phi(t/\sigma)$. ∎
10.2. Asymptotic normality of spectral estimators
Before stating the main CLT, we establish two technical lemmas that make the NSA proof rigorous.
Lemma 10.3 (Infinitesimal perturbation preserves standard-part distribution).
Let $Y,Z$ be finite hyperreal random variables on $({}^{*}\Omega,L({}^{*}\mathsf{P}))$ such that $Y\approx Z$ holds $L({}^{*}\mathsf{P})$-almost surely. Then $\operatorname{st}(Y)$ and $\operatorname{st}(Z)$ have the same distribution under $L({}^{*}\mathsf{P})$.
Proof.
Let $A:=\{\omega:Y(\omega)\approx Z(\omega)\}$. By hypothesis, $L({}^{*}\mathsf{P})(A)=1$. For any $t\in\mathbb{R}$,
$\{\operatorname{st}(Y)\leqslant t\}\cap A=\{\operatorname{st}(Z)\leqslant t\}\cap A$
because if $Y\approx Z$ and $\operatorname{st}(Y)\leqslant t$, then $\operatorname{st}(Z)=\operatorname{st}(Y)\leqslant t$, and vice versa. Since $L({}^{*}\mathsf{P})(A^{c})=0$, we have
$L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t)=L({}^{*}\mathsf{P})(\operatorname{st}(Y)\leqslant t,A)=L({}^{*}\mathsf{P})(\operatorname{st}(Z)\leqslant t,A)=L({}^{*}\mathsf{P})(\operatorname{st}(Z)\leqslant t).\qed$
Lemma 10.4 (Asymptotic linearity in first-order form).
Under the hypotheses of Theorem 10.5, the asymptotic linearity of the spectral L-statistic can be expressed as follows: for every $\varepsilon\in\mathbb{Q}_{>0}$, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geqslant n_{0}$,
(26) $\mathsf{P}\Big(\Big|\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathrm{IF}_{\varphi}(X_{i})\Big|>\varepsilon\Big)<\varepsilon,$
where $\mathrm{IF}_{\varphi}(x):=\int_{0}^{1}\varphi(\alpha)\,q_{X}^{\prime}(\alpha)\,(\mathbf{1}_{x\leqslant q_{X}(\alpha)}-\alpha)\,d\alpha$. This statement is first-order in the sense that it can be formalised using only bounded quantification over rationals, hence is amenable to transfer and underspill.
Proof.
This is the content of [17, Thm. 8.5]: under the smoothness hypotheses (bounded $\varphi$, $q_{X}\in C^{1}$ with finite $\int\varphi^{2}(q_{X}^{\prime})^{2}$), the L-statistic $\hat{\rho}_{n,\varphi}$ is asymptotically linear with influence function $\mathrm{IF}_{\varphi}$. The $o_{\mathsf{P}}(1)$” remainder term means precisely that for every $\varepsilon>0$, $\mathsf{P}(|\text{remainder}|>\varepsilon)\to 0$, which can be restated in the $\forall\varepsilon\exists n_{0}\forall n\geqslant n_{0}$” form above. ∎
Theorem 10.5 (Asymptotic normality of spectral plug-in CREs).
Let $X\in L^{2}$ and $\varphi$ a bounded spectrum. Assume the quantile function $q_{X}$ is continuously differentiable on $(0,1)$ with derivative $q_{X}^{\prime}=1/f_{X}(q_{X})$, where $f_{X}$ is the density of $X$. Assume moreover that $\int_{0}^{1}\varphi(\alpha)^{2}q_{X}^{\prime}(\alpha)^{2}\,d\alpha<\infty$, so that $\sigma_{\varphi}^{2}$ in (28) is finite. Then
(27) $\sqrt{n}\left(\hat{\rho}_{n,\varphi}(X_{1},\ldots,X_{n})-\rho_{\varphi}(X)\right)\xrightarrow{d}N(0,\sigma_{\varphi}^{2}),$
where
(28) $\sigma_{\varphi}^{2}=\int_{0}^{1}\int_{0}^{1}(\min(\alpha,\beta)-\alpha\beta)\varphi(\alpha)\varphi(\beta)q_{X}^{\prime}(\alpha)q_{X}^{\prime}(\beta)\,d\alpha\,d\beta.$
Proof.
The result is a classical asymptotic normality theorem for L-statistics; see [17, Ch. 8] or [18, Ch. 21] for the functional delta method approach. For the reader’s convenience, we provide a complete NSA proof using the hyperfinite CLT and the lemmata above.
The influence function representation. For L-statistics of the form $\hat{\rho}_{n,\varphi}=-\int_{0}^{1}q_{n}(\alpha)\varphi(\alpha)\,d\alpha$, the classical theory provides the asymptotic linearity result stated in Lemma 10.4. The influence function is
$\mathrm{IF}_{\varphi}(x):=\int_{0}^{1}\varphi(\alpha)\,q_{X}^{\prime}(\alpha)\,\bigl(\mathbf{1}_{x\leqslant q_{X}(\alpha)}-\alpha\bigr)\,d\alpha.$
Under the smoothness assumptions, $\mathrm{IF}_{\varphi}\in L^{2}$ with $\mathsf{E}[\mathrm{IF}_{\varphi}(X)]=0$ and $\operatorname{\mathsf{Var}}(\mathrm{IF}_{\varphi}(X))=\sigma_{\varphi}^{2}$.
NSA formulation via hyperfinite CLT. For infinite $N\in{}^{*}\mathbb{N}$, consider the internal average
$T_{N}:=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\mathrm{IF}_{\varphi}(X_{i}).$
By the hyperfinite CLT (Theorem 10.1), $\operatorname{st}(T_{N})\sim N(0,\sigma_{\varphi}^{2})$.
By Lemma 10.4, the asymptotic linearity statement (26) holds for all $n\geqslant n_{0}$ for some standard $n_{0}$. For the fixed infinite $N$ in our standing model, $N>n_{0}$, so
$^{*}\mathsf{P}\Big(\Big|\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))-T_{N}\Big|>\varepsilon\Big)<\varepsilon.$
Intersecting over a sequence $\varepsilon=1/m$ ($m\in\mathbb{N}$) and taking Loeb probabilities, we see that $\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))\approx T_{N}$ holds $L({}^{*}\mathsf{P})$-almost surely.
Conclusion via standard part. By Lemma 10.3, since $\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))\approx T_{N}$ $L({}^{*}\mathsf{P})$-a.s., we have
$\operatorname{st}\bigl(\sqrt{N}(\hat{\rho}_{N,\varphi}-\rho_{\varphi}(X))\bigr)\sim\operatorname{st}(T_{N})\sim N(0,\sigma_{\varphi}^{2}).$
Transfer back to standard sequences. To obtain the standard convergence in distribution, we use overspill on a complement set. Fix $t\in\mathbb{Q}$ and rational $\varepsilon>0$. Define the internal set
$S_{t,\varepsilon}:=\{n\in{}^{*}\mathbb{N}:|{}^{*}\mathsf{P}(\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))\leqslant t)-\Phi(t/\sigma_{\varphi})|<\varepsilon\}.$
This set is internal because the defining predicate uses $^{*}\mathsf{P}$ (the non-standard extension of $\mathsf{P}$) and involves only internal random variables, hypernatural $n$, and standard constants $t,\varepsilon,\sigma_{\varphi},\Phi$. Choose a standard $m\in\mathbb{N}$ so large that $1/m<\varepsilon/3$ and
$\Phi\!\left(\frac{t+1/m}{\sigma_{\varphi}}\right)-\Phi\!\left(\frac{t-1/m}{\sigma_{\varphi}}\right)<\varepsilon/3,$
which is possible by continuity of $\Phi$.
Let us define the influence function of the spectral functional by
$\mathrm{IF}_{\varphi}(x):=\int_{0}^{1}\varphi(\alpha)\,q_{X}^{\prime}(\alpha)\bigl(\mathbf{1}_{\{x\leqslant q_{X}(\alpha)\}}-\alpha\bigr)\,d\alpha=\int_{0}^{1}\varphi(\alpha)\,\frac{\mathbf{1}_{\{x\leqslant q_{X}(\alpha)\}}-\alpha}{f_{X}(q_{X}(\alpha))}\,d\alpha.$
For $n\in{}^{*}\mathbb{N}$ we define the (internal) hyperfinite sum
$T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathrm{IF}_{\varphi}(X_{i}).$
By transfer of (26) with $\varepsilon=1/m$, every infinite $n$ satisfies
$^{*}\mathsf{P}\big(|\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))-T_{n}|>1/m\big)<1/m.$
Moreover, the classical CLT for the i.i.d. sequence $\mathrm{IF}_{\varphi}(X_{i})$ implies (by transfer) that for each standard $u\in\mathbb{R}$ and each infinite $n$,
$^{*}\mathsf{P}(T_{n}\leqslant u)\approx\Phi(u/\sigma_{\varphi}).$
For such $n$ we therefore have the sandwich bounds
$^{*}\mathsf{P}(T_{n}\leqslant t-1/m)-1/m\leqslant{}^{*}\mathsf{P}(\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))\leqslant t)\leqslant{}^{*}\mathsf{P}(T_{n}\leqslant t+1/m)+1/m.$
Consequently, for every infinite $n$,
$\Big|{}^{*}\mathsf{P}(\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))\leqslant t)-\Phi(t/\sigma_{\varphi})\Big|$ $\leqslant\max_{\pm}\Big|{}^{*}\mathsf{P}(T_{n}\leqslant t\pm 1/m)-\Phi\!\Big(\frac{t\pm 1/m}{\sigma_{\varphi}}\Big)\Big|$
$\quad+\Big|\Phi\!\Big(\frac{t+1/m}{\sigma_{\varphi}}\Big)-\Phi\!\Big(\frac{t-1/m}{\sigma_{\varphi}}\Big)\Big|+\frac{1}{m}<\varepsilon,$
so all infinite $n$ lie in $S_{t,\varepsilon}$.
Let $B_{t,\varepsilon}:={}^{*}\mathbb{N}\setminus S_{t,\varepsilon}$ be the internal complement. Since $N\in S_{t,\varepsilon}$ for every infinite $N$, the set $B_{t,\varepsilon}$ contains no infinite hypernaturals. If $B_{t,\varepsilon}$ contained arbitrarily large standard integers, then by overspill (Proposition 3.21) it would contain an infinite hypernatural, a contradiction. Hence there exists $n_{0}\in\mathbb{N}$ such that $B_{t,\varepsilon}\subseteq\{1,\dots,n_{0}-1\}$, and therefore every standard $n\geqslant n_{0}$ lies in $S_{t,\varepsilon}$.
Since $\varepsilon$ and $t$ were arbitrary rationals, and CDFs are determined by their values at rationals (by right-continuity), we obtain $\sqrt{n}(\hat{\rho}_{n,\varphi}-\rho_{\varphi}(X))\xrightarrow{d}N(0,\sigma_{\varphi}^{2})$. ∎
11. Extensions to Orlicz hearts
We briefly discuss how the hyperfinite framework extends to coherent risk measures on Orlicz hearts.
11.1. Orlicz spaces and duality
Let $\Phi:[0,\infty)\to[0,\infty)$ be a Young function (convex, increasing, $\Phi(0)=0$, $\Phi(x)/x\to\infty$). The Orlicz space $L^{\Phi}$ consists of random variables $X$ with $\mathsf{E}[\Phi(|X|/\lambda)]<\infty$ for some $\lambda>0$. The Orlicz heart $H^{\Phi}\subseteq L^{\Phi}$ consists of $X$ with this condition holding for all $\lambda>0$.
The dual Orlicz space $L^{\Psi}$, where $\Psi$ is the complementary Young function, provides the dual pairs for robust representation.
11.2. Hyperfinite representation on Orlicz hearts
Remark 11.1 (Orlicz hearts: what changes).
A full Orlicz-heart analogue of Theorem 4.3 requires replacing the $L^{\infty}$$L^{1}$ duality by the $H^{\Phi}$$L^{\Psi}$ duality (where $\Psi$ is complementary to $\Phi$) and imposing the usual lower semicontinuity/Fatou-type condition in the $H^{\Phi}$ topology; see [6].
Specifically, the key modifications are:
  1. (i)
    The representing set $\mathcal{Z}$ in Theorem 4.3 must consist of densities $Z\in L^{\Psi}_{+}$ with $\mathsf{E}[Z]=1$, and compactness is taken in $\sigma(L^{\Psi},H^{\Phi})$.
  2. (ii)
    The internal lifting $\tilde{X}$ of $X\in H^{\Phi}$ must satisfy $\frac{1}{N}\sum_{k=1}^{N}\Phi(|\tilde{X}(k)|/\lambda)$ finite for all standard $\lambda>0$, not merely boundedness.
  3. (iii)
    For $Z^{\sharp}\in{}^{*}\mathcal{Z}$, $S$-integrability of the product $\tilde{X}\cdot Z^{\sharp}$ follows from the Orlicz–Hölder inequality $|\mathsf{E}[XZ]|\leqslant 2\|X\|_{\Phi}\|Z\|_{\Psi}$ (transferred internally).
  4. (iv)
    The normalisation $H(Z)\approx 1$ still holds since $\mathsf{E}[Z]=1$ transfers.
The remainder of the proof of Theorem 4.3 proceeds unchanged once these modifications are in place.
11.3. Consistency on Orlicz domains
For plug-in consistency on $H^{\Phi}$, the uniform boundedness assumption (K3) in Theorem 8.1 is replaced by:
  1. (K3’)
    There exists $C<\infty$ such that $|\widehat{\operatorname{ES}}_{\alpha,n}(x)|\leqslant C\|x\|_{\Phi}$ for all $\alpha,n,x$.
Under this and the other assumptions (with $L^{\infty}$ replaced by $H^{\Phi}$), the Kusuoka plug-in consistency theorem extends to Orlicz hearts, though a complete treatment requires careful attention to the Orlicz-space duality theory.
12. Concluding remarks
We have developed a systematic hyperfinite framework for coherent risk estimation, demonstrating that non-standard analysis provides both conceptual clarity and technical power for this class of problems.
The hyperfinite viewpoint suggests several directions for future work:
  • Sensitivity analysis. Internal Lipschitz properties of hyperfinite risk functionals could yield robustness bounds for CREs under model misspecification.
  • High-dimensional extensions. The hyperfinite framework may extend to systemic risk measures on high-dimensional portfolios, where the number of assets grows with the sample size.
  • Dynamic risk. Time-consistent dynamic risk measures might admit hyperfinite backward induction representations, simplifying certain asymptotic analyses.
  • Computational aspects. The hyperfinite picture suggests natural discretisation schemes for computing coherent risk measures, with explicit error bounds derived from the standard-part construction.
Acknowledgements
The author thanks the Scientific Circle of Financial Mathematics at Jagiellonian University (Koło Naukowe Matematyki Finansowej UJ) for the invitation to speak at the 26th edition of the “Future Financier Academy” (Akademia Przyszłego Finansisty) seminar held in Kacwin, 28–30 November 2025. It was at this conference that the author first learned of the arXiv preprint [1] by Aichele, Cialenco, Jelito, and Pitera on coherent estimation of risk measures, and where the idea of reformulating their results using non-standard analysis was first conceived.